### v1

# Random Turán Problems for Hypergraph Expansions

Abstract
The random Turán number $\mathrm{ex}(G_{n,p}^{r},F)$ is the maximum number of edges in an $F$-free subgraph of the random $r$-uniform hypergraph $G_{n,p}^{r}$. We prove general results which (informally) shows that if $F$ is an $r_{0}$-graph, then upper bounds for $\mathrm{ex}(G_{n,p}^{r_{0}},F)$ can be lifted into upper bounds for $\mathrm{ex}(G_{n,p}^{r},F^{(r)})$ where $F^{(r)}$ is the $r$-uniform expansion of $F$, i.e. the hypergraph obtained from $F$ by inserting $r-r_{0}$ distinct vertices into each edge of $F$. These results unify and generalize most known upper bounds for random Turán numbers of degenerate hypergraphs of uniformity at least 3, and also provide new tight bounds for the random Turán numbers of expansions of theta graphs and complete bipartite graphs.
1 Introduction
Given an $r$-uniform hypergraph $F$ (or $r$-graph for short), we recall that the Turán number $\mathrm{ex}(n,F)$ is the maximum number of edges that an $n$-vertex $F$-free $r$-graph can have. Let $G_{n,p}^{r}$ denote the random $r$-graph on $n$ vertices obtained by including each possible edge independently and with probability $p$, and when $r=2$ we simply write $G_{n,p}$ instead of $G_{n,p}^{2}$. We define the random Turán number $\mathrm{ex}(G_{n,p}^{r},F)$ to be the maximum number of edges in an $F$-free subgraph of $G_{n,p}^{r}$. Note that when $p=1$ we have $\mathrm{ex}(G_{n,1}^{r},F)=\mathrm{ex}(n,F)$, so the random Turán number can be viewed as a probabilistic analog of the classical Turán number.
The asymptotics for $\mathrm{ex}(G_{n,p}^{r},F)$ are essentially known if $F$ is not an $r$-partite $r$-graph due to independent breakthrough work of Conlon and Gowers [3] and of Schacht [21], and because of this we focus only on the degenerate case when $F$ is $r$-partite. This degenerate case seems to be very difficult even in the graph setting $r=2$ where effective upper bounds are known only for even cycles [10, 5], complete bipartite graphs [10], and theta graphs [8]; and effective lower bounds are known only for powers of rooted trees [22]. All of these bounds agree with a recent conjecture of McKinely and Spiro [8], which gives a general prediction for how the function $\mathrm{ex}(G_{n,p},F)$ should behave for all bipartite $F$.
For $r$-graphs with $r\geq 3$, the only $F$ for which effective bounds are known are complete $r$-partite $r$-graphs [23], loose even cycles [13, 15], Berge cycles [15, 23], expansions of hypergraph cliques and subgraphs of tight trees [18, 14], and non-Sidorenko hypergraphs [17]. We note that part of the difficulty for hypergraphs lies in the fact that no analog of the McKinley-Spiro conjecture mentioned above is known to hold for $r$-graphs, which is closely related to the lack of a hypergraph analog for Sidorenko’s conjecture [17].
The primary aim of this paper is to expand on the current knowledge of random Turán numbers for hypergraphs by studying the following wide class of examples.
Definition 1.
Given an $r_{0}$-graph $F$, we define its $r$-expansion $F^{(r)}$ to be the $r$-graph obtained by inserting $r-r_{0}$ distinct new vertices into each edge of $F$. When $F=C_{\ell}$ is a graph cycle we will sometimes refer to $F^{(r)}=C_{\ell}^{(r)}$ as a loose cycle.
Hypergraph expansions are a natural and well-studied class of hypergraphs, especially in the context of (classical) Turán numbers. For example, the famed Erdős matching conjecture [4] is equivalent to determining $\mathrm{ex}(n,M_{k}^{(r)})$ when $M_{k}$ is a graph matching of size $k$. There are many more results for Turán numbers of expansions [2, 6, 11, 19], and we refer the interested reader to the survey by Mubayi and Verstraëte [12] for much more on (classical) Turán problems for expansions.
Hyeprgraph expansions are also well-studied (though not in a systematic way) in the context of random Turán numbers, see for example [13, 18, 14, 15] mentioned above. This is in large part because, as was initially observed by Mubayi and Yepremyan [13], one can sometimes [^3] Provided these bounds come from a balanced supersaturation result; see Section 1.2 for more. take known bounds for the random Turán problem of $F$ and “lift” these to bounds for the random Turán problem of expansions $F^{(r)}$.
Somewhat more precisely, given an $r$-graph $H$, we say that a set of vertices $S$ with $|S|=r_{0}$ is an $r_{0}$-shadow of $H$ if there exists an edge of $H$ containing the set $S$. Mubayi and Yepremyan thus obtained their upper bounds for the random Turán number of loose even cycles by (roughly speaking) looking at the 2-shadows of $G^{r}_{n,p}$ and then applying known random Turán bounds for graph even cycles due to Morris and Saxton [10] as a black box. A similar (though ultimately inequivalent) approach was used by Nie [15] by inductively looking at the $(r-1)$-shadows of $G^{r}_{n,p}$ rather than the 2-shadows. In total, these two approaches combine to give the following bounds, where here we say that a sequence of events $A_{n}$ holds asymptotically almost surely or a.a.s. for short if $\Pr[A_{n}]\to 1$ as $n\to\infty$, and we write $f(n)\ll g(n)$ if $f(n)/g(n)\to 0$ as $n\to\infty$.
Theorem 1.1 ([15, 13]).
For every $r\geq 4$ and $\ell\geq 2$, we have a.a.s.
$\mathrm{ex}(G^{r}_{n,p},C^{(r)}_{2\ell})=\left\{\begin{aligned} &\max\{\Theta(pn^{r-1}),n^{1+\frac{1}{2\ell-1}}(\log n)^{\Theta(1)}\},~{}~{}&\text{if}~{}p\gg n^{-r+1+\frac{1}{2\ell-1}}\\ &(1+o(1))p\binom{n}{r},~{}~{}&\text{if}~{}n^{-r}\ll p\ll n^{-r+1+\frac{1}{2\ell-1}}.\\ \end{aligned}\right.$
We emphasize here the peculiar situation that tight bounds for loose cycles are not known when $r=3$ despite the tight bounds known for $r=2$ [10] and for $r\geq 4$ due to 1.1.
1.1 Main Results: Random Turán Problems
The aim of this paper is to give a systematic generalization of the approaches of Mubayi-Yepremyan and Nie for loose even cycles mentioned above in order to prove a series of “Lifting Theorems” which informally allows one to lift bounds for the random Turán problem from hypergraphs $F$ into bounds for expansions $F^{(r)}$. Our Lifting Theorems are technical to state, so we postpone their formal statements until the next subsection and instead discuss some of their applications here.
Our Lifting Theorems together with known results on the random Turán problem for graph even cycles $C_{2\ell}$ can be used to prove 1.1, and more generally, these theorems together with known results can be used to recover every previously known upper bound for random Turán problems of expansions [13, 18, 14, 15]. These Lifting Theorems can also be used to obtain many new results as well. For example, while tight bounds for $C_{4}^{(3)}=K_{2,2}^{(3)}$ remain unknown, we can prove tight bounds for all other $K_{2,t}$ expansions.
Theorem 1.2.
If $r,t\geq 2$ are integers such that either $r,t\geq 3$ or $r\geq 4$, then for all $0<p=p(n)\leq 1$ we have a.a.s.
$\mathrm{ex}(G^{r}_{n,p},K_{2,t}^{(r)})=\left\{\begin{aligned} &\max\{\Theta(pn^{r-1}),n^{\frac{3t-1}{2t-1}}(\log n)^{\Theta(1)}\},~{}~{}&\text{if}~{}p\gg n^{-r+\frac{3t-1}{2t-1}}\\ &(1+o(1))p\binom{n}{r},~{}~{}&\text{if}~{}n^{-r}\ll p\ll n^{-r+\frac{3t-1}{2t-1}}.\\ \end{aligned}\right.$
More generally, our results extend to all $K_{s,t}$ at the cost of a suboptimal bound on $r$.
Theorem 1.3.
If $r,s,t\geq 2$ are integers such that $t\geq s$ and either $t\geq s^{2}-2s+3$ and $r\geq s+1$ or $r\geq s+2$, then for all $0<p=p(n)\leq 1$ we have a.a.s.
$\mathrm{ex}(G^{r}_{n,p},K_{s,t}^{(r)})=\left\{\begin{aligned} &\max\{\Theta(pn^{r-1}),n^{2+\frac{s+t-2}{st-1}}(\log n)^{\Theta(1)}\},~{}~{}&\text{if}~{}p\gg n^{-r+2+\frac{s+t-2}{st-1}}\\ &(1+o(1))p\binom{n}{r},~{}~{}&\text{if}~{}n^{-r}\ll p\ll n^{-r+2+\frac{s+t-2}{st-1}}.\\ \end{aligned}\right.$
It is possible that these tight bounds for $\mathrm{ex}(G_{n,p}^{r},K_{s,t}^{(r)})$ continue to hold for any $r$ such that (${r\choose 2}\geq s$ as this is precisely the range where $\mathrm{ex}(n,K_{s,t}^{(r)})=\Theta(n^{r-1})$ for $t$ sufficiently large [7, Theorem 1.4], but 1.3 is the best bound for $r$ that can be obtained with our Lifting Theorems and the known (tight) random Turán bounds for the graph $K_{s,t}$ [10]. Nevertheless, it is possible to go beyond 1.3 by applying our Lifting Theorems to the hypergraphs $K_{s,t}^{(3)}$ instead of the graphs $K_{s,t}$. We will do precisely this in forthcoming work [16] where we use our Lifting Theorems together with a somewhat involved supersaturation argument to obtain the tight bounds of 1.3 for $r$ at least roughly $32s/$ 3 for $s\geq 3$.
Finally, our Lifting Theorems gives bounds for expansions of theta graphs. Here the theta graph $\theta_{a,b}$ is defined to be the graph consisting of $a$ internally disjoint paths of length $b$ between a common pair of vertices $u,v$. For example, $\theta_{2,b}=C_{2b}$ and $\theta_{a,2}=K_{2,a}$. Our Lifting Theorems gives good bounds on $\theta_{a,b}^{(r)}$ whenever effective bounds for $\theta_{a,b}$ are known and $a\geq 3$ (see 5.2), and in particular, recent results of McKinley and Spiro [8] give the following.
Theorem 1.4.
If $r\geq 3$, $a\geq 100$, and $b\geq 3$ are integers, then for all $0<p=p(n)\leq 1$, we have a.a.s.
$\mathrm{ex}(G^{r}_{n,p},\theta_{a,b}^{(r)})=\left\{\begin{aligned} &\max\{\Theta(pn^{r-1}),n^{1+\frac{a-1}{ab-1}}(\log n)^{\Theta(1)}\},~{}~{}&\text{if}~{}p\gg n^{-r+1+\frac{a-1}{ab-1}}\\ &(1+o(1))p\binom{n}{r},~{}~{}&\text{if}~{}n^{-r}\ll p\ll n^{-r+1+\frac{a-1}{ab-1}}.\\ \end{aligned}\right.$
1.2 Main Results: Balanced Supersaturation
There are two key ingredients in proving upper bounds for the random Turán problem. The first is hypergraph containers, which is a general method established independently by Balogh, Morris, and Samotij [1] and by Saxton and Thomason [20] for upper bounding the number of independent sets in hypergraphs. The second is balanced supersaturation, which is a concept originally introduced by Morris and Saxton [10].
Informally, a balanced supersaturation result for an $r$-graph $F$ says that if an $r$-graph $H$ has many edges, then one can find a collection $\mathcal{H}$ of copies of $F$ in $H$ which is “spread out” in the sense that no set of edges in $H$ lies in too many copies of $F$ in $\mathcal{H}$. More precisely we have the following definition, where here for a hypergraph $\mathcal{H}$ and $S\subseteq V(\mathcal{H})$ we define the degree $\deg_{\mathcal{H}}(S)$ to be the number of edges of $\mathcal{H}$ containing $S$, and we define $\Delta_{i}(\mathcal{H})=\max\{\deg_{\mathcal{H}}(S):S\subseteq V(\mathcal{H}),\ |S|=i\}$.
Definition 2.
Given positive functions $M=M(n),\ \gamma=\gamma(n)$, and $\tau=\tau(n,m)$, we say that an $r$-graph $F$ is $(M,\gamma,\tau)$-balanced if for every $n$-vertex $m$-edge $r$-graph $H$ with $n$ sufficiently large and $m\geq M(n)$, there exists a non-empty collection $\mathcal{H}$ of copies of $F$ in $H$ (which we view as an $|F|$-uniform hypergraph on $E(H)$) such that, for all integers $1\leq i\leq|F|$,
$\Delta_{i}(\mathcal{H})\leq\frac{\gamma(n)|\mathcal{H}|}{m}\left(\frac{\tau(n,m)}{m}\right)^{i-1}.$ (1)
In other words, $F$ is $(M,\gamma,\tau)$-balanced if for every $r$-graph with sufficiently many vertices and a large number of edges (i.e. $m\geq M(n)$), we can find a collection of copies of $F$ which are spread out in the sense of (1).
This definition together with a standard application of hypergraph containers yields the following result. We note that particular cases of this result have appeared throughout earlier works on this subject, and we expect that this general version of the result will be quite convenient to use in future study of random Turán problems (see also 4.1 for general lower bounds).
Proposition 1.5.
Let $F$ be an $r$-graph with $r\geq 2$. If there exists a $C>0$ and positive functions $M=M(n)$ and $\tau=\tau(n,m)$ such that
  • (a)
    $F$ is $(M,(\log n)^{C},\tau)$-balanced, and
  • (b)
    For all sufficiently large $n$ and $m\geq M(n)$, the function $\tau(n,m)$ is non-increasing with respect to $m$ and satisfies $\tau(n,m)\geq 1$,
then there exists $C^{\prime}\geq 0$ such that for all sufficiently large $n$, $m\geq M(n)$, and $0<p\leq 1$ with $pm\rightarrow\infty$ as $n\rightarrow\infty$, we have a.a.s.
$\mathrm{ex}(G^{r}_{n,p},F)\leq\max\left\{C^{\prime}pm,\tau(n,m)(\log n)^{C^{\prime}}\right\}.$
We emphasize that condition (b) here is quite mild: for any $\tau$ such that (a) holds, we can always apply 1.5 after replacing $\tau$ with the new function $\tau^{\prime}(n,m):=1+\max_{m^{\prime}\geq m}\tau(n,m^{\prime})$ since (a) continues to hold if we increase $\tau$.
With 1.5 in mind, our goal now will be to prove a set of Lifting Theorems which show that expansions $F^{(r)}$ are balanced whenever $F$ is. Our first result of this form can be seen as a generalization of the proof of [13, Theorem 5.1, Part 1]. While the statement is rather technical, we emphasize that it essentially just says that if $F$ is $(M,\gamma,\tau)$-balanced for some reasonable $M,\gamma,\tau$, then every expansion $F^{(r)}$ is also balanced in some appropriate way.
Theorem 1.6.
Let $r>r_{0}\geq 2$ and let $F$ be an $r_{0}$-graph such that there exist positive functions $M_{r_{0}}=M_{r_{0}}(n)$, $\gamma_{r_{0}}=\gamma_{r_{0}}(n)$ and $\tau_{r_{0}}=\tau_{r_{0}}(n,m)$ with the following properties:
  • (a)
    $F$ is $(M_{r_{0}},\gamma_{r_{0}},\tau_{r_{0}})$-balanced,
  • (b)
    For sufficiently large $n$, the function $M_{r_{0}}(n)n^{-\frac{r-r_{0}}{r-1}}$ is non-decreasing with respect to $n$,
  • (c)
    For sufficiently large $n$, the function $\gamma_{r_{0}}(n)$ is non-decreasing with respect to $n$.
  • (d)
    For sufficiently large $n$, $m\geq M_{r_{0}}(n)$, and any real $x\geq 1$, the function $\tau_{r_{0}}(n,m)$ is non-increasing with respect to $m$ and $\tau_{r_{0}}(nx,mx^{\frac{r-r_{0}}{r-1}})$ is non-decreasing with respect to $x$.
Then there exists a sufficiently large constant $C$ depending only on $F$ and $r$ such that for the functions
$M_{r}(n):=\max\left\{M_{r_{0}}(n)^{\frac{r-1}{r_{0}-1}}n^{-\frac{r-r_{0}}{r_{0}-1}},~{}n^{r-1}\right\}(\log n)^{C},$
$\gamma_{r}(n):=\gamma_{r_{0}}(n)(\log n)^{C},$
and
$\tau_{r}(n,m):=\tau_{r_{0}}\left(n,n^{\frac{r-r_{0}}{r-1}}m^{\frac{r_{0}-1}{r-1}}(\log n)^{-C}\right)(\log n)^{C},$
we have that $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced.
Our second Lifting Theorem, which can be viewed as an extension of the proof of Theorem 2.6 and Theorem 2.7 in [15], gives stronger bounds than 1.6 when going from $F^{(r-1)}$ to $F^{(r)}$ but can only be applied when $r$ is sufficiently large. To state this, we require a few extra definitions.
An $r$-graph $T$ is said to be a tight $r$-tree if its edges can be ordered as $h_{1},\dots,~{}h_{t}$ so that
$\forall i\geq 2~{}\exists v\in h_{i}~{}and~{}1\leq s\leq i-1~{}such~{}that~{}v\not\in\cup_{j=1}^{i-1}h_{j}~{}and~{}h_{i}-v\subset h_{s}.$
Given an $r$-graph $F$ with at least 2 edges, we denote its $r$-density by
$d_{r}(F):=\max_{F^{\prime}\subseteq F,\ |F^{\prime}|\geq 2}\left\{\frac{|F^{\prime}|-1}{v(F^{\prime})-r}\right\}.$
Theorem 1.7.
Let $F$ be an $(r-1)$-graph with $r-1,|F|\geq 2$ such that $F^{(r)}$ is a spanning subgraph of a tight $r$-tree and such that there exist positive functions $M_{r-1}=M_{r-1}(n)$, $\gamma_{r-1}=\gamma_{r-1}(n)$ and $\tau_{r-1}=\tau_{r-1}(n,m)$ with the following properties:
  • (a)
    $F$ is $(M_{r-1},\gamma_{r-1},\tau_{r-1})$-balanced,
  • (b)
    For sufficiently large $n$ the function $\gamma_{r-1}(n)$ is non-decreasing with respect to $n$,
  • (c)
    For sufficiently large $n$ and $m\geq M_{r-1}(n)$, the function $\tau_{r-1}(n,m)$ is non-decreasing with respect to $n$ and non-increasing with respect to $m$.
Then there exists a sufficiently large constant $C$ depending only on $F$ such that the following holds. Let
$M_{r}(n):=Cn^{r-1},$
$\gamma_{r}(n):=\gamma_{r-1}(n)\cdot C\log n,$
and let $A=A(n,m)$ be any function such that, for all sufficiently large $n$ and $m\geq M_{r}(n)$,
  • (d)
    $mn^{1-r}\leq A(n,m)\leq\frac{m}{M_{r-1}(n)\cdot C\log n},$
and let
$\tau_{r}(n,m):=\max\left\{A(n,m)^{-\frac{1}{d_{r}\left(F^{(r)}\right)}}\cdot m,~{}\tau_{r-1}\left(n,~{}\frac{m}{A(n,m)\cdot C\log n}\right)\cdot C\log n\right\}.$
Then $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced.
1.7 turns out to be essentially at least as strong as 1.6 whenever it applies; see the arXiv only appendix for more on this. On the other hand, 1.7 is incapable of proving bounds on $F$ such as $K_{s,t}^{(r)}$ with $r\leq s$ since the tight tree condition does not hold here . As such, our forthcoming work [16] on improving the bound on $r$ in 1.3 will require us to work exclusively with 1.6.
By using 1.7, we will show that if $F$ is an $r_{0}$-graph which has “optimal” balanced supersaturation (see 3.1) and Turán number $O(n^{r_{0}-1})$ (i.e. one whose extremal construction has the same size as that of an $r_{0}$-uniform star), then every expansion of $F$ has essentially optimal bounds for the random Turán problem. For minor technical reasons we restrict our result to $F$ with maximum degree $\Delta(F)$ not too large or small.
Theorem 1.8.
Let $F$ be an $r_{0}$-graph with $2\leq\Delta(F)<|F|$ which is the spanning subgraph of a tight $r_{0}$-tree. If there exists a constant $C_{r_{0}}$ such that $F$ is $(M_{r_{0}},\gamma_{r_{0}},\tau_{r_{0}})$-balanced where $M_{r_{0}}(n)=C_{r_{0}}n^{r_{0}-1}$, $\gamma_{r_{0}}(n)=(\log n)^{C_{r_{0}}}$ and $\tau_{r_{0}}(n,m)=n^{r_{0}-\frac{1}{d_{r_{0}}(F)}}(\log n)^{C_{r_{0}}}$; then for all $r\geq r_{0}$, we have
$\mathrm{ex}(G^{r}_{n,p},F^{(r)})=\left\{\begin{aligned} &\max\{\Theta(pn^{r-1}),n^{r-\frac{1}{d_{r}(F^{(r)})}}(\log n)^{\Theta(1)}\},~{}~{}&\text{if}~{}p\gg n^{-\frac{1}{d_{r}(F^{(r)})}}\\ &(1+o(1))p\binom{n}{r},~{}~{}&\text{if}~{}n^{-r}\ll p\ll n^{-\frac{1}{d_{r}(F^{(r)})}}.\\ \end{aligned}\right.$
[…]
Organization. The rest of our paper is organized as follows. We prove Theorems 1.6 and 1.7 in Section 2, which is the bulk of the technical work of this paper. We then prove several results related to optimal balanced supersaturation in Section 3 and general random Turán results in Section 4, culminating in a proof of 1.8. Finally, we apply our Lifting Theorems to give short proofs of Theorems 1.2 and 1.4 in Section 5, with the proof of 1.3 being deferred to an arXiv only appendix.
2 Balanced Supersaturation for Expansions
2.1 Proof Sketch
In this section we prove our two general results Theorems 1.6 and 1.7 about lifting balanced supersturation from $F$ to the expansions $F^{(r)}$. To illustrate our methods, we informally consider the toy problem of finding copies of $C_{4}^{(3)}$ in a 3-graph $H$ where every pair of vertices in $H$ has degree either 0 or $d$. In this setting we use one of two approaches to build our copies of $C_{4}^{(3)}$.
The first approach grows a $C_{4}^{(3)}$ from a copy of $C_{4}$ in the shadow graph $\partial H$, say a $C_{4}$ which has vertex set $v_{1},v_{2},v_{3},v_{4}$. Because $\{v_{1},v_{2}\}$ has degree $d$ in $H$, there exist at least $d-4$ choices for a vertex $u_{1,2}\notin\{v_{1},v_{2},v_{3},v_{4}\}$ such that $\{v_{1},v_{2},u_{1,2}\}$ is an edge in $H$. Similarly there are at least $d-8$ choices for each of the vertices $u_{2,3},u_{3,4},u_{1,4}$ such that $\{v_{i},v_{j},u_{i,j}\}$ are edges in $H$ and such that all of these vertices are distinct, and these vertices in total define a copy of $C_{4}^{(3)}$ in $H$. With this procedure, each graph $C_{4}$ in $\partial H$ can be used to build roughly $d^{4}$ copies of $C_{4}^{(3)}$ in $H$. As such, if we can find a collection $\mathcal{H}^{\prime}$ of copies of $C_{4}$ in $\partial H$, then we can use this to find a collection $\mathcal{H}$ of copies of $C_{4}^{(3)}$ in $H$ of size roughly $d^{4}|\mathcal{H}^{\prime}|$, and moreover, it will turn out that $\Delta_{i}(\mathcal{H})$ will be small provided $\Delta_{i}(\mathcal{H}^{\prime})$ is, from which we will conclude that balanced supersaturation for $C_{4}$ leads to balanced supersaturation for $C_{4}^{(3)}$.
The second approach involves growing a $C_{4}^{(3)}$ from an arbitrary edge $h=\{v_{1},v_{2},v_{3}\}$ in $H$. Because $v_{1},v_{3}$ has degree $d$, we can choose in $d-3$ ways some $v_{4}\notin\{v_{1},v_{2},v_{3}\}$ which is in an edge with $v_{1},v_{3}$, after which we can perform the exact same expansion procedure as above to build a total of roughly $d^{5}$ copies of $C_{4}^{(3)}$ from this edge $h$. In total this approach gives a collection of $C_{4}^{(3)}$’s of size about $d^{5}|H|$, which will do better than the previous approach depending on the relative sizes of $d$ and $|H|$. We emphasize here that while our first approach works for essentially any $F$ and $r$, this second approach will only work if the uniformity of $H$ is sufficiently large for the tight tree condition to hold.
We informally refer to the first approach described here as “Shadow Expansion” and the second as “Greedy Expansion”. We will prove 1.6 by using only shadow expansion, and we will prove 1.7 by combining both of these methods.
2.2 Shadow Expansion: Proof of 1.6
We begin with our main lemma for the Shadow Expansion procedure described above. Recall that given an $r$-graph $H$, we say that a set of vertices $S$ with $|S|=r_{0}$ is an $r_{0}$-shadow of $H$ if there exists an edge of $H$ containing the set $S$.
Lemma 2.1.
Let $r>r_{0}\geq 2$. Let $F$ be an $r_{0}$-graph such that there exist functions $M=M(n)$, $\gamma=\gamma(n)$ and $\tau=\tau(n,m)$ with the property that
  • (A)
    $F$ is $(M,\gamma,\tau)$-balanced.
Let $H$ be an $n$-vertex $r$-partite $r$-graph with $r$-partition $V(H)=V_{1}\cup\dots\cup V_{r}$. Let $\hat{n}=|V_{1}\cup\dots\cup V_{r_{0}}|$ and let $\hat{m}$ be the number of $r_{0}$-shadows of $H$ in $V_{1}\cup\dots\cup V_{r_{0}}$. If
  • (B)
    $\hat{n}$ is sufficiently large,
  • (C)
    $\hat{m}\geq M(\hat{n})$, and if
  • (D)
    there exist real numbers $d$ and $D$ with $D\geq d\geq 2v(F^{(r)})n^{r-r_{0}-1}$ such that for every $r_{0}$-shadow $S$ of $H$ in $V_{1}\cup\dots\cup V_{r_{0}}$,
    $d\leq\deg_{H}(S)\leq D,$
then there exists a collection $\mathcal{H}$ of copies of $F^{(r)}$ in $H$ such that for all $1\leq i\leq|F|$,
$\Delta_{i}(\mathcal{H})\leq\frac{\gamma(\hat{n})|\mathcal{H}|}{D\hat{m}}\cdot\left(\frac{\tau(\hat{n},\hat{m})}{D\hat{m}}\right)^{i-1}\cdot\left(\frac{2^{r+1}D}{d}\right)^{|F|}.$
Proof.
Let $H_{[r_{0}]}$ be the $r_{0}$-graph on $V_{1}\cup\dots\cup V_{r_{0}}$ whose edges are all of the $r_{0}$-shadows of $H$ in $V_{1}\cup\dots\cup V_{r_{0}}$. Note that $v(H_{[r_{0}]})=\hat{n}$ and $|H_{[r_{0}]}|=\hat{m}$. By properties (a), (b) and (c), there exists a collection $\hat{\mathcal{H}}$ of $F$ in $H_{[r_{0}]}$ such that for all $1\leq i\leq|F|$,
$\Delta_{i}(\hat{\mathcal{H}})\leq\frac{\gamma(\hat{n})|\hat{\mathcal{H}}|}{\hat{m}}\cdot\left(\frac{\tau(\hat{n},\hat{m})}{\hat{m}}\right)^{i-1}.$ (2)
We now obtain a collection $\mathcal{H}$ of copies of $F^{(r)}$ in $H$ by “expanding” each copy of $F$ in $\hat{\mathcal{H}}$. More precisely, consider the following procedure for constructing copies of $F^{(r)}$ in $H$. Start by picking some $\hat{F}\in\hat{\mathcal{H}}$ and give an arbitrary ordering $e_{1},\ldots,e_{m}$ to the edges of $\hat{F}$. We then iteratively for each $1\leq k\leq m$ pick $h_{k}\in H$ to be any edge of $H$ which contains $e_{k}$ and which is disjoint from $(V(\hat{F})\cup\bigcup_{k^{\prime}<k}h_{k^{\prime}})\setminus\{e_{k}\}$, after which the process terminates. Observe that if this process terminates, then these $h_{k}$ edges form a copy of $F^{(r)}$ in $H$, and we let $\mathcal{H}$ denote the set of copies of $F^{(r)}$ that can be formed in this way. It remains to analyze the properties of this collection.
Observe that in the procedure, the number of choices for $h_{k}$ given the edge $e_{k}$ is crudely at least
$\deg_{H}(e_{k})-v(F^{(r)})n^{r-r_{0}-1},$ (3)
since $|(V(\hat{F})\cup\bigcup_{k^{\prime}<k}h_{k^{\prime}})\setminus\{e_{k}\}|\leq v(F^{(r)})$ and since there are at most $n^{r-r_{0}-1}$ edges of $H$ containing $e_{k}$ and any given vertex from $(V(\hat{F})\cup\bigcup_{k^{\prime}<k}h_{k^{\prime}})\setminus\{e_{k}\}$. Since (3) is at least $d/2$ by property (d), we find that the total number of ways of going through the procedure is at least |^$|\hat{\mathcal{H}}|(d/2)^{|F|}$. Moreover, since each copy of $F^{(r)}$ can be produced in at most $2^{r|F|}$ ways from this procedure (since there are crudely at most (${r\choose r_{0}}^{|F|}$ ways of picking an $\hat{F}$ which produces $F^{(r_{0})}$), we in total find that
|||^$|\mathcal{H}|\geq|\hat{\mathcal{H}}|\left(\frac{d}{2^{r+1}}\right)^{|F|}.$ (4)
Now fix some set $\sigma\subseteq H$ of $i$ edges of $H$ and observe that one can specify each copy of $F^{(r)}$ in $\mathcal{H}$ containing $\sigma$ through the following two steps:
  • 1.
    Specify a copy $\hat{F}\in\hat{\mathcal{H}}$ which contains all $i$ of the $[r_{0}]$-shadows of $\sigma$ in $\bigcup_{i\in[r_{0}]}V_{i}$. There are at most $\Delta_{i}(\hat{\mathcal{H}})$ ways to do this.
  • 2.
    For each of the $|F|-i$ other $[r_{0}]$-shadows of $\hat{F}$ that are not in $\sigma$, we specify an edge in $H$ which contains it. By property (d), for each $[r_{0}]$-shadow there are at most $D$ ways to do this.
In total then, we see for all $1\leq i\leq|F|$ that
$\Delta_{i}(\mathcal{H})\leq\Delta_{i}(\hat{\mathcal{H}})D^{|F|-i}.$ (5)
Finally, by (5), (2), and (4), we have, for all $1\leq i\leq|F|$,
$\Delta_{i}(\mathcal{H})$ $\leq\frac{\gamma(\hat{n})|\hat{\mathcal{H}}|}{\hat{m}}\cdot\left(\frac{\tau(\hat{n},\hat{m})}{\hat{m}}\right)^{i-1}\cdot D^{|F|-i}$
$\leq\frac{\gamma(\hat{n})|\mathcal{H}|}{\hat{m}}\cdot\left(\frac{\tau(\hat{n},\hat{m})}{\hat{m}}\right)^{i-1}\cdot D^{|F|-i}\cdot\left(\frac{d}{2^{r+1}}\right)^{-|F|}$
$=\frac{\gamma(\hat{n})|\mathcal{H}|}{D\hat{m}}\cdot\left(\frac{\tau(\hat{n},\hat{m})}{D\hat{m}}\right)^{i-1}\cdot\left(\frac{2^{r+1}D}{d}\right)^{|F|}.$
In order to use Lemma 2.2, we will need to work with a hypergraph whose codegrees $\deg_{H}(\sigma)$ are well controlled. We do this through the following regularization lemma.
Lemma 2.2.
If $r\geq 2$ and $H$ is an $n$-vertex $r$-partite $r$-graph on $V_{1}\cup\cdots\cup V_{r}$ with $n$ sufficiently large in terms of $r$, then there exists a subgraph $H^{\prime}\subseteq H$ without isolated vertex and a real-valued vector $\Delta$ indexed by $2^{[r]}$ such that $|H^{\prime}|\geq(r\log n)^{-2^{r}}|H|/2$ and for any set $I\subseteq[r]$, every $|I|$-shadow $S$ of $H^{\prime}$ in $\bigcup_{i\in I}V_{i}$ satisfies
$\Delta_{I}(2r\log n)^{-2^{r}}\leq\deg_{H^{\prime}}(S)\leq\Delta_{I}.$
Moreover, let $n^{\prime}=|V(H^{\prime})|$ and let $\partial_{I}$ denote the number of $|I|$-shadows of $H^{\prime}$ in $\bigcup_{i\in I}V_{i}$. It is possible to relabel the indices of $V_{1},\dots,V_{r}$ such that for any $1\leq k\leq r$,
$\partial_{[k]}\geq(n^{\prime})^{\frac{r-k}{r-1}}|H^{\prime}|^{\frac{k-1}{r-1}}(2r\log n)^{-k2^{r}},$
and
$\Delta_{[k]}\leq\left(\frac{|H^{\prime}|}{n^{\prime}}\right)^{\frac{r-k}{r-1}}\left(2r(\log n)^{k+1}\right)^{2^{r}}.$
Proof.
Given an integral vector $y$ indexed by $2^{[r]}$, we say that an edge $e\in H$ is of type $y$ if for every set $I\subseteq[r]$, we have $2^{y_{I}-1}\leq\deg_{H}(e\cap\bigcup_{i\in I}V_{i})<2^{y_{I}}$. By the pigeonhole principle, there exists some $y\in[\log(n^{r})]^{2^{r}}$ such that at least $\log(n^{r})^{-2^{r}}|H|$ edges of $H$ are of type $y$. Let $H_{0}$ be the collection of edges of type $y$, and let $\Delta$ be the vector with $\Delta_{I}:=2^{y_{I}}$.
Iteratively, given $H_{0}$, if there exists some set $I\subseteq[r]$ and $S\subseteq\bigcup_{i\in I}V_{i}$ with $|S|=|I|$ such that $0<\deg_{H_{0}}(S)<2^{-2^{r}}\log(n^{r})^{-2^{r}}\Delta_{I}$, then delete every edge from $H_{0}$ containing $S$. Let $H^{\prime}$ be the resulting hypergraph, noting that it satisfies the desired degree conditions by construction (since in particular, $\deg_{H^{\prime}}(S)\leq\deg_{H_{0}}(S)\leq\Delta_{I}$ for any $S\subseteq\bigcup_{i\in I}V_{i}$). Further, note that for any set $I\subseteq[r]$, the number of sets $S\subseteq\bigcup_{i\in I}V_{i}$ of size $|I|$ with positive degree in $H_{0}$ is at most $2|H|/\Delta_{I}$ (since any set with positive degree in $H_{0}$ has degree at least $\Delta_{I}/2$ in $H$ by definition of $H_{0}$), so the total number of edges deleted from $H_{0}$ to obtain $H^{\prime}$ is at most
$2^{-2^{r}}\log(n^{r})^{-2^{r}}\Delta_{I}\cdot 2|H|/\Delta_{I}=2^{-2^{r}+1}\log(n^{r})^{-2^{r}}|H|.$
Summing this bound over all $I\subseteq[r]$ gives
$|H^{\prime}|\geq|H_{0}|-2^{r}\cdot 2^{-2^{r}+1}\log(n^{r})^{-2^{r}}|H|\geq\log(n^{r})^{-2^{r}}|H|-\frac{1}{2}\log(n^{r})^{-2^{r}}|H|\geq\frac{1}{2}\log(n^{r})^{-2^{r}}|H|,$
giving the desired lower bound on $|H^{\prime}|$. It thus remains to prove the “moreover” part of the statement. For this we make use of the inequalities
$\partial_{I}\Delta_{I}(2r\log n)^{-2^{r}}\leq|H^{\prime}|\leq\partial_{I}\Delta_{I},$ (6)
which follow from $|H^{\prime}|=\sum_{S\subseteq\bigcup_{i\in I}V_{i}}\deg_{H^{\prime}}(S)$ for any $I$ and our bounds for these degrees .
We may assume without loss of generality that $H^{\prime}$ has no isolated vertices, as deleting these do not change any of the previously mentioned properties. We will show by induction on $k$ that, upon relabeling indices of $V_{i}$, we have for all $1\leq k\leq r$ that
$\partial_{[k]}\geq(n^{\prime})^{\frac{r-k}{r-1}}|H^{\prime}|^{\frac{k-1}{r-1}}(2r\log n)^{-k2^{r}}.$
By the Pigeonhole Principle, we can assume without loss of generality that $\partial_{\{1\}}=|V(H^{\prime})\cap V_{1}|\geq n^{\prime}/r$. Let $12\leq k\leq r-$ 1 and suppose we have inductively shown
$\partial_{[k-1]}\geq(n^{\prime})^{\frac{r-k+1}{r-1}}|H^{\prime}|^{\frac{k-2}{r-1}}(2r\log n)^{-(k-1)2^{r}}.$ (7)
Assume for contradiction that $\partial_{[k-1]\cup\{j\}}<(n^{\prime})^{\frac{r-k}{r-1}}|H^{\prime}|^{\frac{k-1}{r-1}}(2r\log n)^{-k2^{r}}$ for every $k\leq j\leq r$. In particular, this and (6) implies
$\Delta_{[k]}\geq\frac{|H^{\prime}|}{\partial_{[k]}}\geq\left(\frac{|H^{\prime}|}{n^{\prime}}\right)^{\frac{r-k}{r-1}}(2r\log n)^{k2^{r}}.$ (8)
Claim 2.3.
For every $k$-shadow $\{v_{1},\ldots,v_{k}\}$ of $H^{\prime}$ with $v_{i}\in V_{i}$ for all $i$, the number of vertices $u$ such that $\{v_{1},\ldots,v_{k},u\}$ is a $(k+1)$-shadow of $H^{\prime}$ is at least
$\left(\frac{\Delta_{[k]}}{(2r\log n)^{2^{r}}}\right)^{\frac{1}{r-k}}.$
Proof.
Let $U$ denote the set of vertices $u$ as described in the claim. Observe that every edge containing $\{v_{1},\ldots,v_{k}\}$ in $H^{\prime}$ consists of itself together with $r-k$ vertices of $U$, hence
$\deg_{H^{\prime}}(v_{1},\ldots,v_{k})\leq{|U|\choose r-k}\leq|U|^{r-k},$
and rearranging gives
$|U|\geq\deg_{H^{\prime}}(v_{1},\ldots,v_{k})^{\frac{1}{r-k}}\geq\left(\frac{\Delta_{[k]}}{(2r\log n)^{2^{r}}}\right)^{\frac{1}{r-k}},$
with this last step using our lower bound for $I$-degrees that we verified in the first half of the proof. ∎
Observe that if $\{v_{1},\ldots,v_{k}\}$ is a $k$-shadow of $H^{\prime}$ as in the claim and $\{v_{1},\ldots,v_{k},u\}$ is a $(k+1)$-shadow, then $\{v_{1},\ldots,v_{k-1},u\}$ is a $k$-shadow of $H^{\prime}$ contained in $V_{1}\cup\cdots\cup V_{k-1}\cup V_{j}$ for some $k+1\leq j\leq r$. Hence by the claim together with (7) and (8) we have
$\sum^{r}_{j=k+1}\partial_{[k-1]\cup\{j\}}$ $\geq\partial_{[k-1]}\left(\frac{\Delta_{[k]}}{(2r\log n)^{2^{r}}}\right)^{\frac{1}{r-k}}\geq(n^{\prime})^{\frac{r-k}{r-1}}|H^{\prime}|^{\frac{k-1}{r-1}}(2r\log n)^{-(k-1)2^{r}-\frac{1}{r-k}2^{r}}$
$\geq(r-k)(n^{\prime})^{\frac{r-k}{r-1}}|H^{\prime}|^{\frac{k-1}{r-1}}(2r\log n)^{-k2^{r}},$
where this last inequality holds with equality if $k=r-1$ and otherwise holds for $n$ sufficiently large (since the exponent for the logarithmic factor is strictly more than $-k2^{r}$ in this case).
By the Pigeonhole Principle, we conclude that for some $k+1\leq j\leq r$,
$\partial_{[k-1]\cup\{j\}}\geq(n^{\prime})^{\frac{r-k}{r-1}}|H^{\prime}|^{\frac{k-1}{r-1}}(\log n)^{-k2^{r}}.$
This contradicts our assumption that $\partial_{[k-1]\cup\{j\}}<n^{\prime\frac{r-k}{r-1}}|H^{\prime}|^{\frac{k-1}{r-1}}(2r\log n)^{-k2^{r}}$ for each $k\leq j\leq r$. Hence, by relabeling indices if necessary, we have
$\partial_{[k]}\geq(n^{\prime})^{\frac{r-k}{r-1}}|H^{\prime}|^{\frac{k-1}{r-1}}(2r\log n)^{-k2^{r}},$
completing our inductive lower bound for $\partial_{[k]}$. Finally, by (6) we have for all $1\leq k\leq r$ that
$\Delta_{[k]}\leq\frac{(2r\log n)^{2^{r}}|H^{\prime}|}{\partial_{[k]}}\leq\left(\frac{|H^{\prime}|}{n^{\prime}}\right)^{\frac{r-k}{r-1}}\left(2r(\log n)^{k+1}\right)^{2^{r}}.$
We now have all the tools we need to prove our first main result 1.6, which we restate below for convenience.
Theorem 1.6.
Let $r>r_{0}\geq 2$ and let $F$ be an $r_{0}$-graph such that there exist positive functions $M_{r_{0}}=M_{r_{0}}(n)$, $\gamma_{r_{0}}=\gamma_{r_{0}}(n)$ and $\tau_{r_{0}}=\tau_{r_{0}}(n,m)$ with the following properties:
  • (a)
    $F$ is $(M_{r_{0}},\gamma_{r_{0}},\tau_{r_{0}})$-balanced,
  • (b)
    For sufficiently large $n$, the function $M_{r_{0}}(n)n^{-\frac{r-r_{0}}{r-1}}$ is non-decreasing with respect to $n$,
  • (c)
    For sufficiently large $n$, the function $\gamma_{r_{0}}(n)$ is non-decreasing with respect to $n$.
  • (d)
    For sufficiently large $n$, $m\geq M_{r_{0}}(n)$, and any real $x\geq 1$, the function $\tau_{r_{0}}(n,m)$ is non-increasing with respect to $m$ and $\tau_{r_{0}}(nx,mx^{\frac{r-r_{0}}{r-1}})$ is non-decreasing with respect to $x$.
Then there exists a sufficiently large constant $C$ depending only on $F$ and $r$ such that for the functions
$M_{r}(n):=\max\left\{M_{r_{0}}(n)^{\frac{r-1}{r_{0}-1}}n^{-\frac{r-r_{0}}{r_{0}-1}},~{}n^{r-1}\right\}(\log n)^{C},$
$\gamma_{r}(n):=\gamma_{r_{0}}(n)(\log n)^{C},$
and
$\tau_{r}(n,m):=\tau_{r_{0}}\left(n,n^{\frac{r-r_{0}}{r-1}}m^{\frac{r_{0}-1}{r-1}}(\log n)^{-C}\right)(\log n)^{C},$
we have that $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced.
Proof.
Let $H$ be an $n$-vertex $m$-edge $r$-graph where $n$ is sufficiently large and $m\geq M_{r}(n)$. Let $H^{\prime}$ be a maximum $r$-partite subgraph of $H$, noting that $|H^{\prime}|\geq\frac{r\,!}{r^{r}}\cdot m$ by considering a random $r$-partition of $H$. By Lemma 2.2 there exists a vector $\Delta$ and a subgraph $H^{\prime\prime}$ of $H^{\prime}$ on $n^{\prime\prime}$ vertices with $r$-partition $V_{1}\cup\dots\cup V_{r}$ such that
$|H^{\prime\prime}|\geq 2^{-1}(r\log n)^{-2^{r}}|H^{\prime}|\geq\frac{r\,!\,m}{2r^{r}(r\log n)^{2^{r}}}\geq\frac{m}{(\log n)^{2^{r}+1}},$ (9)
with this last bound holding for $n$ sufficiently large; and moreover, every $r_{0}$-shadow $S$ of $H^{\prime\prime}$ in $\bigcup_{i\in[r_{0}]}V_{i}$ satisfies
$(2r\log n)^{-2^{r}}\Delta_{[r_{0}]}\leq\deg_{H^{\prime\prime}}(S)\leq\Delta_{[r_{0}]},$ (10)
where
$\Delta_{[r_{0}]}\leq\left(\frac{|H^{\prime\prime}|}{n^{\prime\prime}}\right)^{\frac{r-r_{0}}{r-1}}\left(2r(\log n)^{r_{0}+1}\right)^{2^{r}}\leq\left(\frac{|H^{\prime\prime}|}{n^{\prime\prime}}\right)^{\frac{r-r_{0}}{r-1}}(\log n)^{r_{0}2^{r+1}}.$ (11)
Let $H_{[r_{0}]}$ be the $r_{0}$-graph on $\bigcup_{i\in[r_{0}]}V_{i}$ whose edges are all $r_{0}$-shadows of edges of $H^{\prime\prime}$ in $\bigcup_{i\in[r_{0}]}V_{i}$. Let $\hat{n}=|\bigcup_{i\in[r_{0}]}V_{i}|$ and let $\hat{m}=|H_{[r_{0}]}|$. Our aim now will be to apply 2.1 to $H_{[r_{0}]}$, and to this end we need to effectively lower bound $\hat{m}$ and, in view of (10), we will also need to show that $d:=\Delta_{[r_{0}]}(2r\log n)^{-2^{r}}$ is at least $2v(F^{(r)})n^{r-r_{0}-1}$. Towards both of these goals, we observe by (10) that
$|H^{\prime\prime}|\leq\Delta_{[r_{0}]}\hat{m}.$ (12)
By (12) and (9), we have
$\Delta_{[r_{0}]}\geq\frac{|H^{\prime\prime}|}{\hat{m}}\geq\frac{m}{\hat{m}(\log n)^{2^{r}+1}}.$
Further, using the trivial bounds $\hat{m}\leq n^{r_{0}}$ and $m\geq M_{r}(n)\geq n^{r-1}(\log n)^{2^{r+1}+2}$ for $C$ sufficiently large, we have
$(2r\log n)^{-2^{r}}\Delta_{[r_{0}]}\geq 2v(F^{(r)})n^{r-r_{0}-1}.$ (13)
Turning to $\hat{m}$, we have by (12), (11), and (9) that
$\hat{m}\geq\frac{|H^{\prime\prime}|}{\Delta_{[r_{0}]}}\geq(n^{\prime\prime})^{\frac{r-r_{0}}{r-1}}|H^{\prime\prime}|^{\frac{r_{0}-1}{r-1}}(\log n)^{-r_{0}2^{r+1}}\geq(n^{\prime\prime})^{\frac{r-r_{0}}{r-1}}\left(\frac{m}{(\log n)^{2^{r}+1}}\right)^{\frac{r_{0}-1}{r-1}}(\log n)^{-r_{0}2^{r}}.$ (14)
Using (14) together with $m\geq M_{r}(n)$ and property (b), we have for $C$ sufficiently large that
$\hat{m}$ $\geq(n^{\prime\prime})^{\frac{r-r_{0}}{r-1}}\left(\frac{M_{r_{0}}(n)^{\frac{r-1}{r_{0}-1}}n^{-\frac{r-r_{0}}{r_{0}-1}}(\log n)^{C}}{(\log n)^{2^{r}+1}}\right)^{\frac{r_{0}-1}{r-1}}(\log n)^{-r_{0}2^{r}}$
$\geq(n^{\prime\prime})^{\frac{r-r_{0}}{r-1}}\left(M_{r_{0}}(n^{\prime\prime})^{\frac{r-1}{r_{0}-1}}(n^{\prime\prime})^{-\frac{r-r_{0}}{r_{0}-1}}\right)^{\frac{r_{0}-1}{r-1}}$
$=M_{r_{0}}(n^{\prime\prime})$
$\geq M_{r_{0}}(\hat{n}).$
Observe that we may assume $\hat{n}$ is sufficiently large. Indeed, we assume in the statement that $n$ is sufficiently large, which combined with $m\geq M_{r}(n)\geq n^{r-1}$ and (14) implies $\hat{m}=|H_{[r_{0}]}|$ is sufficiently large, from which it follows that $\hat{n}=v(H_{[r_{0}]})$ is sufficiently large. Using this and property (a), we may apply Lemma 2.1 to $H^{\prime\prime}$, giving a collection $\mathcal{H}$ of copies of $F^{(r)}$ such that for all $1\leq i\leq|F|$,
$\Delta_{i}(\mathcal{H})\leq\frac{\gamma_{r_{0}}(\hat{n})|\mathcal{H}|}{\Delta_{r_{0}}\hat{m}}\cdot\left(\frac{\tau_{r_{0}}(\hat{n},\hat{m})}{\Delta_{r_{0}}\hat{m}}\right)^{i-1}\left(2^{r+1}(2r\log n)^{2^{r}}\right)^{|F|}.$ (15)
Using this with (9) and (12) gives,
$\Delta_{i}(\mathcal{H})$ $\leq\frac{\left(2^{r+1}(2r\log n)^{2^{r}}\right)^{|F|}\cdot\gamma_{r_{0}}(\hat{n})|\mathcal{H}|}{|H^{\prime\prime}|}\cdot\left(\frac{\tau_{r_{0}}(\hat{n},\hat{m})}{|H^{\prime\prime}|}\right)^{i-1}$ (16)
$\leq\frac{2r^{r}(r\log n)^{2^{r}}}{r\,!}\cdot\frac{\left(2^{r+1}(2r\log n)^{2^{r}}\right)^{|F|}\cdot\gamma_{r_{0}}(n)|\mathcal{H}|}{m}\cdot\left(\frac{2r^{r}(r\log n)^{2^{r}}}{r\,!}\cdot\frac{\tau_{r_{0}}(\hat{n},\hat{m})}{m}\right)^{i-1}.$
Note that, by both parts of property (d) and by (14),
$\tau_{r_{0}}(\hat{n},\hat{m})$ $\leq\tau_{r_{0}}\left(\hat{n},\hat{n}^{\frac{r-r_{0}}{r-1}}\left(\frac{r\,!\,m}{2r^{r}(r\log n)^{2^{r}}}\right)^{\frac{r_{0}-1}{r-1}}(\log n)^{-r_{0}2^{r}}\right)$ (17)
$\leq\tau_{r_{0}}\left(n,n^{\frac{r-r_{0}}{r-1}}\left(\frac{r\,!\,m}{2r^{r}(r\log n)^{2^{r}}}\right)^{\frac{r_{0}-1}{r-1}}(\log n)^{-r_{0}2^{r}}\right).$
Finally, by property (c), the definition of $\gamma_{r}$ and $\tau_{r}$, and (16) and (17), we conclude that for all $1\leq i\leq|F|$,
$\Delta_{i}(\mathcal{H})\leq\frac{\gamma_{r}(n)|\mathcal{H}|}{m}\cdot\left(\frac{\tau_{r}(n,m)}{m}\right)^{i-1}.$
Therefore, $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced for $n$ sufficiently large by definition. ∎
2.3 Greedy Expansion: Proof of 1.7
We now prove our second result 1.7 which lifts balancedness from $(r-1)$-graphs $F$ to $F^{(r)}$ whenever $F^{(r)}$ is contained in a tight $r$-tree. We begin by introducing two lemmas that will be useful. The following says that every $r$-partite $r$-graph contains a large subgraph such that either all $(r-1)$-shadows have large codegrees or all $(r-1)$-shadows in some $(r-1)$ parts have small (but not too small) and regular codegrees.
Lemma 2.4.
For $r\geq 2$, let $H$ be an $r$-partite $r$-graph on $n$ vertices, then for any $|H|/(4n^{r-1})<A\leq n$, one of the following two statements is true:
  • (i)
    There exists a subgraph $\hat{H}\subseteq H$ with $|\hat{H}|\geq|H|/2$ such that every $(r-1)$-shadow of $\hat{H}$ has codegree at least $A$.
  • (ii)
    There exist a subgraph $\hat{H}\subseteq H$ with $|\hat{H}|\geq|H|/(4r\log n)$, an $r$-partition $V(\hat{H})=V_{1}\cup\dots\cup V_{r}$, and
    $\frac{|H|}{4n^{r-1}}<D\leq A$
    such that any $(r-1)$-shadow $\sigma$ of $\hat{H}$ in $V_{1}\cup\dots\cup V_{r-1}$ satisfies
    $D/2\leq\deg_{\hat{H}}(\sigma)<D.$
Proof.
Let $V(H)=V_{1}\cup\dots\cup V_{r}$ be an $r$-partition of $H$. We run the following algorithm that partitions $H$ into subgraphs $\hat{H}_{L}$ and $\hat{H}_{J,a}$, where $J\in\binom{[r]}{r-1}$ and $1\leq a\leq\log A$:
  • 1.
    Let $H_{0}=H$. Given $H_{i}$ for some $i\geq 0$, if there is an $(r-1)$-shadow $\sigma$ of $H_{i}$ such that $\deg_{H_{i}}(\sigma)\leq A$, then we arbitrarily fix such a $\sigma$ and let $H_{i+1}=H_{i}-N_{H_{i}}(\sigma)$, where $N_{H_{i}}(\sigma)$ denotes the set of edges in $H_{i}$ containing $\sigma$. Let $(J,a)$ be the unique pair such that $\sigma$ contain a vertex of $V_{j}$ for each $j\in J$ and $2^{a-1}\leq\deg_{H_{i}}(\sigma)<2^{a}$. We put all the elements in $N_{H_{i}}(\sigma)$ into the edge set of $\hat{H}_{J,a}$.
  • 2.
    If all $(r-1)$-shadows $\sigma$ of $H_{i}$ have $\deg_{H_{i}}(\sigma)>A$, then we let $\hat{H}_{L}=H_{i}$ and terminate the algorithm.
If $|\hat{H}_{L}|\geq|H|/2$, we let $\hat{H}=\hat{H}_{L}$, which by definition implies $\hat{H}$ satisfies (i). Otherwise,
$\sum_{1\leq a\leq\log A}\sum_{J\in\binom{[r]}{r-1}}|\hat{H}_{J,a}|>|H|/2.$
Note that each $(r-1)$-shadow $\sigma$ of $H$ contributes at most $2^{a}$ to some $|\hat{H}_{J,a}|$, and the number of $(r-1)$-shadows is at most $n^{r-1}$. Thus we have
$\sum_{1\leq a\leq\log\frac{|H|}{4n^{r-1}}}\sum_{J\in\binom{[r]}{r-1}}|\hat{H}_{J,a}|\leq n^{r-1}\frac{|H|}{4n^{r-1}}=\frac{|H|}{4},$
and hence
$\sum_{\log\frac{|H|}{4n^{r-1}}<a\leq\log A}\sum_{J\in\binom{[r]}{r-1}}|\hat{H}_{J,a}|>\frac{|H|}{4}.$
By the Pigeonhole Principle, there exists a pair $(J,a)$ such that $2^{a}\geq\frac{|H|}{4n^{r-1}}$ and that $|\hat{H}_{J,a}|\geq|H|/(4r\log n)$. Fix such a pair $(J,a)$ and let $\hat{H}=\hat{H}_{J,a}$. Without loss of generality, we can suppose $J=\{1,\dots,r-1\}$. By definition, any $(r-1)$-shadow $\sigma$ in $V_{1}\cup\dots\cup V_{r-1}$ has $2^{a-1}\leq\deg_{\hat{H}}(\sigma)<\min\{2^{a},A\}$. Therefore, $\hat{H}$ satisfies (ii) by letting $D=\min\{2^{a},A\}\geq|H|/(4n^{r-1})$. ∎
The following lemma comes from [14, Lemma 2.2] and is proven in essentially the same way as described in the second approach of Section 2.1.
Lemma 2.5 ([14] Lemma 2.2).
Let $T$ be a tight $r$-tree and let $F$ be a spanning subgraph of $T$ with $|F|\geq 2$. There exists a constant $C>0$ such that the following holds for all sufficiently large $n$ and every $A$ with $C\leq A\leq\binom{n}{r}/n^{r-1}$. Let $H$ be an $r$-graph on $n$ vertices. If every $(r-1)$-shadow $\sigma$ of $H$ has $\deg_{H}(\sigma)>A$, then there exists a collection $\mathcal{H}$ of copies of $F$ in $H$ satisfying
  1. (a)
    $|\mathcal{H}|\geq C^{-1}|H|A^{v(F)-r}$;
  2. (b)
    $\Delta_{i}(\mathcal{H})\leq CA^{v(F)-r-\frac{i-1}{d_{r}(F)}}$,  $1\leq i\leq|F|$.
In particular, we have for all $1\leq i\leq|F|$ that
$\Delta_{i}(\mathcal{H})\leq C^{2}\frac{|\mathcal{H}|}{|H|}A^{-\frac{i-1}{d_{r}(F)}}.$
We are now ready to prove 1.7, which we restate below for convenience.
Theorem 1.7.
Let $F$ be an $(r-1)$-graph with $r-1,|F|\geq 2$ such that $F^{(r)}$ is a spanning subgraph of a tight $r$-tree and such that there exist positive functions $M_{r-1}=M_{r-1}(n)$, $\gamma_{r-1}=\gamma_{r-1}(n)$ and $\tau_{r-1}=\tau_{r-1}(n,m)$ with the following properties:
  • (a)
    $F$ is $(M_{r-1},\gamma_{r-1},\tau_{r-1})$-balanced,
  • (b)
    For sufficiently large $n$ the function $\gamma_{r-1}(n)$ is non-decreasing with respect to $n$,
  • (c)
    For sufficiently large $n$ and $m\geq M_{r-1}(n)$, the function $\tau_{r-1}(n,m)$ is non-decreasing with respect to $n$ and non-increasing with respect to $m$.
Then there exists a sufficiently large constant $C$ depending only on $F$ such that the following holds. Let
$M_{r}(n):=Cn^{r-1},$
$\gamma_{r}(n):=\gamma_{r-1}(n)\cdot C\log n,$
and let $A=A(n,m)$ be any function such that, for all sufficiently large $n$ and $m\geq M_{r}(n)$,
  • (d)
    $mn^{1-r}\leq A(n,m)\leq\frac{m}{M_{r-1}(n)\cdot C\log n},$
and let
$\tau_{r}(n,m):=\max\left\{A(n,m)^{-\frac{1}{d_{r}\left(F^{(r)}\right)}}\cdot m,~{}\tau_{r-1}\left(n,~{}\frac{m}{A(n,m)\cdot C\log n}\right)\cdot C\log n\right\}.$
Then $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced.
Proof.
Let $H$ be an $n$-vertex $m$-edge $r$-graph such that $n$ is sufficiently large and $m\geq M_{r}(n)$. Let $H^{\prime}\subseteq H$ be a maximum sized $r$-partite subgraph of $H$. By considering a random $r$-partition of $V(H)$, we know that $|H^{\prime}|\geq\frac{r\,!}{r^{r}}m$. By Lemma 2.4, there exists a subgraph $\hat{H}\subseteq H^{\prime}$ with
$|\hat{H}|\geq\frac{|H^{\prime}|}{4r\log n}\geq\frac{r\,!}{4r^{r+1}}\cdot\frac{m}{\log n}$ (18)
and an $r$-partition $V(\hat{H})=V_{1}\cup\dots\cup V_{r}$ such that one of the following two statements is true.
  • (i)
    Every $(r-1)$-shadow of $\hat{H}$ has codegree at least $A$.
  • (ii)
    There exists $|H^{\prime}|/(4n^{r-1})<D\leq A$ such that every $(r-1)$-shadow $\sigma$ of $\hat{H}$ in $V_{1}\cup\dots\cup V_{r-1}$ satisfies
    $D/2\leq\deg_{\hat{H}}(\sigma)<D.$
The proof splits into two cases accordingly.
Case 1: Statement (i) is true. Because $A\geq mn^{1-r}$ by property (d) and because $m\geq M_{r}(n)=Cn^{r-1}$ with $C$ sufficiently large, we can apply Lemma 2.5 to conclude that there exists a constant $C^{\prime}$ and a collection $\mathcal{H}$ of copies of $F^{(r)}$ in $\hat{H}$ such that for all $1\leq i\leq|F^{(r)}|$,
$\Delta_{i}(\mathcal{H})\leq C^{\prime}\frac{|\mathcal{H}|}{|\hat{H}|}A^{-\frac{i-1}{d_{r}(F^{(r)})}}\leq\frac{C^{\prime}2r^{r+1}\log n}{r\,!}\cdot\frac{|\mathcal{H}|}{m}\left(\frac{A^{-\frac{1}{d_{r}(F^{(r)})}}\cdot m}{m}\right)^{i-1}\leq\gamma_{r}(n)\cdot\frac{|\mathcal{H}|}{m}\left(\frac{\tau_{r}(n,m)}{m}\right)^{i-1},$
where the second inequality used (18), and this last step used the definition of $\gamma_{r}$ and property (b) to conclude
$\gamma_{r}(n)\geq C\log n\cdot\min\{1,\gamma_{r-1}(n_{r-1})\}\geq\frac{C^{\prime}2r^{r+1}\log n}{r!}$
provided $C$ is sufficiently large, and similarly that $\tau_{r}(n,m)\geq A^{-\frac{1}{d_{r}(F^{(r)})}}\cdot m$.
Case 2: Statement (ii) is true. We aim to apply Lemma 2.1 on $\hat{H}$. To this end, let $\hat{H}_{[r-1]}$ be the $(r-1)$-graph on $\cup_{i\in[r-1]}V_{i}$ whose edges are all $(r-1)$-shadows of $\hat{H}$ in $\cup_{i\in[r-1]}V_{i}$. Let $\hat{n}=v(\hat{H}_{[r-1]})=|\cup_{i\in[r-1]}V_{i}|$ and let $\hat{m}=|\hat{H}_{[r-1]}|$.
Recall that $D\geq\frac{|H^{\prime}|}{4n^{r-1}}\geq\frac{r\,!\,m}{4r^{r}n^{r-1}}$ and that $m\geq Cn^{r-1}$. So for $C$ sufficiently large, we have
$D/2\geq\frac{Cr\,!}{4r^{r}}\geq 2v(F^{(r)}).$
By statement (ii), we have
$D\hat{m}\geq|\hat{H}|.$ (19)
Thus by (19), (18), the hypothesis $D\leq A$, and property (d), we have
$\hat{m}\geq\frac{|\hat{H}|}{D}\geq\frac{r\,!}{4r^{r+1}}\cdot\frac{m}{A\cdot\log n}\geq M_{r-1}(n).$
Given sufficiently large $n$, we have sufficiently large $\hat{m}$, and hence sufficiently large $\hat{n}$. So by property (a) and the inequalities above, we can apply Lemma 2.1 on $\hat{H}$ to obtain a collection $\mathcal{H}$ of copies of $F^{(r)}$ such that, for all $1\leq i\leq|F|$,
$\Delta_{i}(\mathcal{H})\leq\frac{\gamma_{r-1}(\hat{n})|\mathcal{H}|}{D\hat{m}}\cdot\left(\frac{\tau_{r-1}(\hat{n},\hat{m})}{D\hat{m}}\right)^{i-1}4^{r|F|}.$
By properties (b), (c) and (d), and inequalities (18) and (19),
$\Delta_{i}(\mathcal{H})$ $\leq\frac{4^{r|F|}2r^{r+1}\log n}{r\,!}\cdot\frac{\gamma_{r-1}(n)|\mathcal{H}|}{m}\cdot\left(\frac{2r^{r+1}\log n}{r\,!}\cdot\frac{\tau_{r-1}\left(n,\frac{r\,!\cdot m}{2r^{r+1}\log n\cdot A}\right)}{m}\right)^{i-1}$
$\leq\frac{\gamma_{r}(n)|\mathcal{H}|}{m}\left(\frac{\tau_{r}(n,m)}{m}\right)^{i-1}.$
In total then, we find that in either case there exists a collection $\mathcal{H}$ satisfying, for all $1\leq i\leq|F|$ that
$\Delta_{i}(\mathcal{H})\leq\frac{\gamma_{r}(n)|\mathcal{H}|}{m}\left(\frac{\tau_{r}(n,m)}{m}\right)^{i-1},$
showing that $F^{(r)}$ is indeed $(M_{r},\gamma_{r},\tau_{r})$-balanced. ∎
3 Optimal Balanced Supersaturation
Here we explore what can be said about (expansions) $F$ which have “optimal” balanced supersaturation. More precisely, we observe the following limits for how balanced a hypergraph $F$ can be.
Proposition 3.1.
If $F$ is an $r$-graph with $r,|F|\geq 2$ which is $(M,\gamma,\tau)$-balanced, then for sufficiently large $n$ we have $M(n)>\mathrm{ex}(n,F)$ and
$\tau(n,m)\geq r^{-\frac{r|F|}{|F|-1}}\gamma(n)^{-\frac{1}{|F|-1}}n^{r-\frac{1}{d_{r}(F)}}.$
Proof.
If $M(n)\leq\mathrm{ex}(n,F)$ for infinitely many $n$, then one could consider an infinite sequence of $F$-free $r$-graphs $H$ with $n$ vertices and $\mathrm{ex}(n,F)$ edges, from which no collection $\mathcal{H}$ of copies of $F$ can be guaranteed, contradicting $F$ being $(M,\gamma,\tau)$-balanced.
Finally, let $H$ be the random $r$-uniform hypergraph $G^{r}_{n,p}$ with $p=m/\binom{n}{r}$ for some $m\geq M(n)$. With positive probability we have $|H|\geq m$ and the number of copies of $F$ in $H$ is at most $p^{|F|}n^{v(F)}\leq r^{r|F|}m^{|F|}n^{v(F)-r|F|}$. Let $\mathcal{H}$ be a non-empty collection of copies of $F$ in $H$ satisfying (1). Then we have $|\mathcal{H}|\leq r^{r|F|}m^{|F|}n^{v(F)-r|F|}$. Note that $\Delta_{|F|}(\mathcal{H})=1$, so by (1) we have
$\tau(n,m)\geq\left(\frac{m}{\gamma(n)|\mathcal{H}|}\right)^{\frac{1}{|F|-1}}m\geq r^{-\frac{r|F|}{|F|-1}}n^{r-\frac{v(F)-r}{|F|-1}}\gamma(n)^{-\frac{1}{|F|-1}}\geq r^{-\frac{r|F|}{|F|-1}}n^{r-\frac{1}{d_{r}(F)}}\gamma(n)^{-\frac{1}{|F|-1}}.$
Our main goal of this section is to prove a result (3.4) giving a sufficient condition for expansions to have optimal balanced supersaturation, which will be the main ingredient for proving 1.8. We will need two lemmas regarding expansions, the first of which shows that $r$-densities play nicely with expansions.
Lemma 3.2.
For any $r_{0}$-graph $F$ with $r_{0},|F|\geq 2$ and integer $r\geq r_{0}$, we have
$\frac{1}{d_{r}(F^{(r)})}=r-r_{0}+\frac{1}{d_{r_{0}}(F)},$
and in particular $d_{r}(F^{(r)})<\frac{1}{r-r_{0}}$.
Proof.
Let $F^{\prime}\subseteq F$ be a subgraph satisfying $d_{r_{0}}(F)=\frac{|F^{\prime}|-1}{v(F^{\prime})-r_{0}}$. Then the expansion $F^{\prime\prime}:=(F^{\prime})^{(r)}\subseteq F^{(r)}$ satisfies
$\frac{1}{d_{r}(F^{(r)})}\leq\frac{v(F^{\prime\prime})-r}{|F^{\prime\prime}|-1}=\frac{v(F^{\prime})+(r-r_{0})|F^{\prime}|-r}{|F^{\prime}|-1}=r-r_{0}+\frac{1}{d_{r_{0}}(F)}.$
Now let $F^{\prime}\subseteq F^{(r)}$ be the subgraph satisfying $d_{r}(F^{(r)})=\frac{|F^{\prime}|-1}{v(F^{\prime})-r}$. Define $F^{\prime\prime}\subseteq F$ to be the induced subraph on $V(F^{\prime})\cap V(F)$. Observe that $|F^{\prime\prime}|\geq|F^{\prime}|$ and $v(F^{\prime\prime})\leq v(F^{\prime})-(r-r_{0})|F^{\prime}|$ (since each edge of $F^{\prime}$ contains $r-r_{0}$ vertices not in $V(F)$). This implies that
$\frac{1}{d_{r_{0}}(F)}\leq\frac{v(F^{\prime\prime})-r_{0}}{|F^{\prime\prime}|-1}\leq\frac{v(F^{\prime})-r_{0}-(r-r_{0})|F^{\prime}|}{|F^{\prime}|-1}=\frac{1}{d_{r}(F^{(r)}}-r+r_{0},$
which together with the previous inequality gives the desired result. ∎
We next have a lemma which gives sufficient conditions for (expansions of) $F$ to be a spanning subgraph of a tight tree. Here the $k$-shadow of an $r_{0}$-graph $T$ is defined to be the $k$-graph on $V(T)$ consisting of the $k$-shadows of $T$.
Lemma 3.3.
If $F$ is a $k$-graph and $r_{0}\geq k$ an integer such that there exists a tight $r_{0}$-tree $T^{\prime}$ whose $k$-shadow contains $F$ as a spanning subgraph, then for all $r\geq r_{0}$ there exists a tight $r$-tree $T$ which contains $F^{(r)}$ as a spanning subgraph.
In particular, if $r_{0}=k$ then the hypothesis of this lemma simply says that $F$ is a spanning subgraph of a tight $r_{0}$-tree.
Proof.
The result is trivial if $F$ is a single edge, so we will assume this is not the case. Given a pair of integers $r\geq r_{0}$ and $\ell\geq k$, we will say that $F$ is $(r,\ell)$-good if there exists some ordering $v_{1},v_{2},\ldots,v_{t}$ of the vertices of $F^{(\ell)}$ and an ordered set of $r$-sets $h_{r},h_{r+1}\ldots,h_{t}$ such that (1) $h_{r}=\{v_{1},\ldots,v_{r}\}$, (2) for all $i>r$ we have $e_{i}\setminus\bigcup_{j<i}h_{j}=\{v_{i}\}$ and there exists some $i^{\prime}<i$ with $h_{i}-v_{i}\subseteq h_{i^{\prime}}$, (3) every edge of $F^{(\ell)}$ is contained in some $h_{i}$.
We first observe that $F$ being $(r,\ell)$-good implies that the $r$-graph $T$ with edges $h_{r},\ldots,h_{t}$ is a tight $r$-tree whose $\ell$-shadow contains $F^{(\ell)}$ as a spanning subgraph. In particular, the hypothesis implies that $F$ is $(r_{0},k)$-good. We aim to show that $F$ is $(r,r)$-good for any $r\geq r_{0}$, proving the result. We do this through double induction.
Assume that we shown $F$ to be $(r,\ell)$-good for some $r\geq r_{0}$ and $r>\ell\geq k$; we aim to show that $F$ is $(r,\ell+1)$-good. Let $h_{r},\ldots,h_{t}$ be the ordered $r$-sets guaranteed by $F$ being $(r,\ell)$-good. For each edge $e\in F^{(\ell)}$, let $v_{e}$ denote the unique vertex such that $\{v_{e}\}\cup e$ is an edge of $F^{(\ell+1)}$, and let $i_{e}$ denote any index such that $h_{i_{e}}$ contains $e$ (which exists by (3)). For each $e\in F^{(\ell)}$, we let $h_{e}$ be any $r$-set which consists of $v_{e}$ and any $(r-1)$-subset of $h_{i_{e}}$ which contains $e$ (which exists since $|e|=\ell<r$ by hypothesis). It is not difficult to check that $h_{r},\ldots,h_{t}$ followed by any ordering of these $h_{e}$ sets verifies that $F$ is $(r,\ell+1)$-good, proving this claim.
Now assume we have shown $F$ to be $(r,r)$-good for some $r\geq r_{0},k$. We aim to show that $F$ is $(r+1,r)$-good. Let $h_{r},\ldots,h_{t}$ be the ordered $r$-sets guaranteed by $F$ being $(r,\ell)$-good. Note that we must have $r<t$, as otherwise (3) would imply that $F^{(r)}$ must in fact just be a single edge, which we assumed not to be the case. With this in mind, we construct $(r+1)$-sets $h_{r+1}^{\prime},\ldots,h_{t}^{\prime}$ by having $h_{i}^{\prime}=h_{i^{\prime}}\cup\{v_{i}\}$ for all $i\geq r+1$. It is not difficult to see that these $(r+1)$-sets verify $F$ as being $(r+1,r)$-good.
Combining the two claims above together with the observation that $F$ is $(r_{0},\ell)$-good gives that $F$ is $(r,r)$-good for all $r\geq r_{0}$, proving the result. ∎
We now prove our main result of the section, which roughly shows that if $F$ has sufficiently high uniformity and optimal balanced supersaturation, then so does every expansion of $F$.
Proposition 3.4.
Let $F$ be an $r_{0}$-graph with $\Delta(F)\geq 2$ which is the spanning subgraph of a tight $r_{0}$-tree such that there exists a constant $C_{r_{0}}$ such that $F$ is $(M_{r_{0}},\gamma_{r_{0}},\tau_{r_{0}})$-balanced with $M_{r_{0}}(n)=C_{r_{0}}n^{r_{0}-1}$, $\gamma_{r_{0}}(n)=(\log n)^{C_{r_{0}}}$ and $\tau_{r_{0}}(n,m)=n^{r_{0}-\frac{1}{d_{r_{0}}(F)}}(\log n)^{C_{r_{0}}}$. Then for all $r\geq r_{0}$, $F^{(r)}$ is the spanning subgraph of a tight tree and there exists a constant $C_{r}$ such that $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced with $M_{r}(n)=C_{r}n^{r-1}$, $\gamma_{r}(n)=(\log n)^{C_{r}}$ and $\tau_{r}(n,m)=n^{r-\frac{1}{d_{r}(F)}}(\log n)^{C_{r}}$.
Proof.
We claim that it suffices to prove the result when $r_{0}=r-1$. Indeed, assuming this case we prove the general result by induction on $r$, the $r=r_{0}$ case being trivial. Assume that we have proven the result up to some $r>r_{0}$. Because the result holds at $r-1$, $F^{(r-1)}$ continues to satisfy the hypothesis of the lemma, so the assumed $r-1$ result completes the induction. With this we may assume that $r_{0}=r-1$ from now on.
It is straightforward to check that properties (a), (b), and (c) of 1.7 are satisfied, and we have that $F^{(r)}$ is a spanning subgraph of a tight tree due to 3.3 applied with $k=r_{0}$. Let $C$ be the constant guaranteed to exist from 1.7. Let $M_{r}(n)=Cn^{r-1}$, let $\gamma_{r}(n)=\gamma_{r-1}(n)\cdot C\log n=C(\log n)^{C_{r-1}+1}$, and let
$A(n,m)=\max\{m^{d_{r}(F^{(r)})}n^{-rd_{r}(F^{(r)})+1},~{}mn^{1-r}\}.$
We claim that property (d) of 1.7 holds. Indeed, the lower bound $A(n,m)\geq mn^{1-r}$ holds trivially since this is one of the terms in the maximum. The upper bound $A(n,m)\leq\frac{m}{M_{r-1}(n)C\log n}=\frac{m}{C_{r-1}n^{r-2}\cdot C\log n}$ holds for the second term of $A$ provided (${n\choose r-1}\geq C_{r-1}n^{r-2}$, which holds for $n$ sufficiently large, and it holds for the first term provided
$m^{d_{r}(F^{(r)})-1}n^{-rd_{r}(F^{(r)})+1}\leq\frac{1}{C_{r-1}\cdot C\log n}n^{2-r}.$
Because $d_{r}(F^{(r)})<1$ by 3.2, this inequality is hardest to satisfy at the smallest possible value of $m\geq Cn^{r-1}$, and one can verify this holds for $n$ sufficiently large through the inequality $-d_{r}(F^{(r)})+2-r<2-r$.
Define
$\tau_{r}(n,m)$ $=\max\left\{A(n,m)^{-\frac{1}{d_{r}\left(F^{(r)}\right)}}\cdot m,~{}\tau_{r-1}\left(n,~{}\frac{m}{A(n,m)\cdot C\log n}\right)\cdot C\log n\right\}$
$\leq n^{r-\frac{1}{d_{r}\left(F^{(r)}\right)}}\cdot C(\log n)^{C_{r-1}+1},$
where this last step used $\tau_{r-1}=n^{r-1-\frac{1}{d_{r-1}(F)}}(\log n)^{C_{r-1}}=n^{r-\frac{1}{d_{r}(F^{(r)}}}(\log n)^{C_{r-1}}$ by 3.2. We conclude by 1.7 that $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced. ∎
4 General Random Turán Results
Here we prove a set of general lower and upper bounds for the random Turán problem, special cases of which have appeared throughout the literature in previous papers. We begin with some simple bounds.
Lemma 4.1.
Let $F$ be an $r$-graph with $\Delta(F)\geq 2$. If $n^{-r}\ll p\ll n^{-\frac{1}{d_{r}(F)}}$, then a.a.s.
$\mathrm{ex}(G_{n,p}^{r},F)=(1+o(1))p{n\choose r}.$
If $p\gg n^{-\frac{1}{d_{r}(F)}}$, then a.a.s.
$\mathrm{ex}(G_{n,p}^{r},F)\geq\max\{\Omega(p\cdot\mathrm{ex}(n,F)),n^{r-\frac{1}{d_{r}(F)}}(\log n)^{-1}\}.$
Proof.
For the upper bound, we observe that trivially $\mathrm{ex}(G_{n,p}^{r},F)\leq|G_{n,p}^{r}|$, and this is at most $(1+o(1))p{n\choose r}$ a.a.s. provided $p\gg n^{-r}$ by the Chernoff bound.
For the first lower bound, let $F^{\prime}\subseteq F$ be such that $d_{r}(F)=\frac{|F^{\prime}|-1}{v(F^{\prime})-r}$, and let $X$ be the number of copies of $F^{\prime}$ in $G_{n,p}^{r}$. Note that $\mathbb{E}[X]=\Theta(p^{|F^{\prime}|}n^{v(F^{\prime})})=o(p{n\choose r})$ by our range of $p$. As such we have $\omega:=\frac{p{n\choose r}}{\mathbb{E}[X]}\to\infty.$ Choose $\epsilon=\epsilon(n)$ any function such that $\epsilon\to 0$ and $\epsilon\omega\to\infty$. By Markov’s inequality,
$\Pr\left[X\geq\epsilon p{n\choose r}\right]\leq\frac{1}{\epsilon\omega}\to 0.$
Thus a.a.s. $X=o(p{n\choose r})$. By the Chernoff bound, $|G_{n,p}^{r}|\geq(1-o(1))p{n\choose r}$ a.a.s. for $p\gg n^{-r}$. Thus by deleting an edge from each of the $X$ copies of $F$ in $G_{n,p}^{r}$, we obtain an $F$-free subgraph of $G_{n,p}^{r}$ with $(1-o(1))p{n\choose r}$ edges a.a.s.
For the remaining lower bounds, we observe that $\Delta(F)\geq 2$ implies (and is in fact equivalent to) $d_{r}(F)>1/r$ by considering $F^{\prime}\subseteq F$ consisting of two intersecting edges. The lower bound $n^{r-\frac{1}{d_{r}(F)}}(\log n)^{-1}$ then follows from the monotonicity of $\mathrm{ex}(G_{n,p}^{r},F)$ with respect to $p$: more precisely, taking $q:=n^{-\frac{1}{d_{r}(F)}}(\log\log n)^{-1}$, we see that $n^{-r}\ll q\ll n^{-\frac{1}{d_{r}(F)}}$ (because $d_{r}(F)>1/r$), and hence the bound above implies that a.a.s. $\mathrm{ex}(G_{n,q}^{r},F)\geq\frac{1}{2}q{n\choose r}\geq n^{r-\frac{1}{d_{r}(F)}}(\log n)^{-1}$, and hence this same lower bound continues to hold a.a.s. for $\mathrm{ex}(G_{n,p}^{r},F)$ for any $p\geq q$.
Finally, the subgraph $G\subseteq G_{n,p}^{r}$ obtained by intersecting $G_{n,p}^{r}$ with an $F$-free subgraph with $\mathrm{ex}(n,F)$ edges is a binomial random variable with mean $p\cdot\mathrm{ex}(n,F)$, so this will be at least $\Omega(p\cdot\mathrm{ex}(n,F))$ a.a.s. provided $p\cdot\mathrm{ex}(n,F)\to\infty$. And indeed, due to the maximum in the lower bound we only need to show this in the case when $p\cdot\mathrm{ex}(n,F)\geq n^{r-\frac{1}{d_{r}(F)}}(\log n)^{-1}\to\infty$ since $d_{r}(F)>1/r$. ∎
We now move onto our upper bound for the random Turán problem 1.5. Our proof uses the hypergaph container method, developed independently by Balogh–Morris–Samotij [1] and Saxton–Thomason [20], which is fundamental in the study of random Turán problems. Here we make use of the following simplified version of Theorem 1.5 in [9]:
Theorem 4.2 ([9]).
For every integer $k\geq 2$, there exists a constant $\epsilon>0$ such that the following holds. Let $B,L\geq 1$ be positive integers and $\mathcal{H}$ a $k$-graph satisfying
$\Delta_{j}(\mathcal{H})\leq\left(\frac{B}{v(\mathcal{H})}\right)^{j-1}\frac{|\mathcal{H}|}{L},~{}~{}\forall 1\leq j\leq k.$ (20)
Then there exists a collection $\mathcal{C}$ of subsets of $V(\mathcal{H})$ such that:
  1. (i)
    For every independent set $I$ of $\mathcal{H}$, there exists $S\in\mathcal{C}$ such that $I\subset S$;
  2. (ii)
    For every $S\in\mathcal{C}$, $|S|\leq v(\mathcal{H})-\epsilon L$;
  3. (iii)
    We have
    $|\mathcal{C}|\leq\exp\left(\frac{\log\left(v(\mathcal{H})\right)B}{\epsilon}\right).$
We now proving 1.5, which we restate below for convenience.
Proposition 1.5.
Let $F$ be an $r$-graph with $r,|F|\geq 2$. If there exist $C>0$ and positive functions $M=M(n)$ and $\tau=\tau(n,m)$ such that
  • (a)
    $F$ is $(M,(\log n)^{C},\tau)$-balanced, and
  • (b)
    For all sufficiently large $n$ and $m\geq M(n)$, the function $\tau(n,m)$ is non-increasing with respect to $m$ and satisfies $\tau(n,m)\geq 1$,
then there exists $C^{\prime}\geq 0$ such that for all sufficiently large $n$, $m\geq M(n)$, and $0<p\leq 1$ with $pm\rightarrow\infty$ as $n\rightarrow\infty$, we have a.a.s.
$\mathrm{ex}(G^{r}_{n,p},F)\leq\max\left\{C^{\prime}pm,\tau(n,m)(\log n)^{C^{\prime}}\right\}.$
Proof.
Let $n$ be sufficiently large and $m\geq M(n)$. Note that by 3.1 we have $m\geq\mathrm{ex}(n,F)=\Omega(n)$ since $r,|F|\geq 2$, and in particular for $n$ sufficiently large we have
$m\geq(\log n)^{C}.$ (21)
We will take $[n]=\{1,\dots,n\}$ to be the vertex set of all the $r$-graphs discussed below. We first prove the following claim by repeatedly applying Theorem 4.2.
Claim 4.3.
There exist a collection $\mathcal{C}$ of $r$-graphs on $[n]$ and a constant $\tilde{C}$ (independent of $n$ and $m$) such that
  • (i)
    Every $F$-free $r$-graph on $[n]$ is contained in some $S\in\mathcal{C}$;
  • (ii)
    $|S|\leq m$ for every $S\in\mathcal{C}$;
  • (iii)
    $|\mathcal{C}|\leq\exp\left((\log n)^{\tilde{C}}\tau(n,m)\right)$.
Proof.
We will iteratively define collections $\mathcal{C}_{i}$ satisfying (i) and the size condition $|\mathcal{C}_{i}|\leq\exp\left(\frac{r\log n\cdot\tau(n,m^{\prime})}{\epsilon}\right)^{i}$ for some $\epsilon$ depending only on $F$ as follows. We start with $\mathcal{C}_{0}=\{K^{r}_{n}\}$, where $K^{r}_{n}$ is the complete $r$-graph on $[n]$ and note that this collection satisfies (i) and $|\mathcal{C}_{0}|=1$. Given a collection $\mathcal{C}_{i-1}$ for $i\geq 1$ which satisfies (i) and the size condition, if $\mathcal{C}_{i-1}$ satisfies (ii) then we stop and set $\mathcal{C}=\mathcal{C}_{i-1}$, and otherwise we construct a new collection $\mathcal{C}_{i}$ as following. Let
$\mathcal{C}_{i-1}^{\leq m}=\{G\in\mathcal{C}_{i-1}:|G|\leq m\}~{}~{}\text{and}~{}~{}\mathcal{C}_{i-1}^{>m}=\{G\in\mathcal{C}_{i-1}:|G|>m\}.$
For each $G\in\mathcal{C}_{i-1}^{>m}$, we construct a collection $\mathcal{C}_{G}$ of $r$-graphs on $[n]$ as follows. Let $m^{\prime}=|G|>m$. By (a), there exists a collection $\mathcal{H}$ of copies of $F$ in $G$ such that, $\forall 1\leq i\leq|F|$,
$\Delta_{i}(\mathcal{H})\leq\frac{(\log n)^{C}|\mathcal{H}|}{m^{\prime}}\left(\frac{\tau(n,m^{\prime})}{m^{\prime}}\right)^{i-1}.$
By applying Theorem 4.2 with $B=\tau(n,m^{\prime})\geq 1$, and $L=m^{\prime}(\log n)^{-C}\geq 1$ (here we use (21) and $m^{\prime}>m$), we obtain a collection $\mathcal{C}_{G}$ of $r$-graphs on $[n]$ (i.e. subsets of (${[n]\choose r}$) such that
  • (i’)
    every $F$-free subgraph of $G$ is contained in some $G^{\prime}\in\mathcal{C}_{G}$;
  • (ii’)
    $|G^{\prime}|\leq\left(1-\epsilon(\log n)^{-C}\right)m^{\prime}$ for every $G^{\prime}\in\mathcal{C}_{G}$;
  • (iii’)
    $|\mathcal{C}_{G}|\leq\exp\left(\frac{r\log n\cdot\tau(n,m^{\prime})}{\epsilon}\right)$.
Here $\epsilon$ is a constant depending only on $|F|$. Having constructed $\mathcal{C}_{G}$ for each $G\in\mathcal{C}_{i-1}^{>m}$, we finally define
$\mathcal{C}_{i}=\mathcal{C}_{i-1}^{\leq m}\cup\bigcup_{G\in\mathcal{C}_{i-1}^{>m}}\mathcal{C}_{G}.$
Note that since we inductively assumed (i) held for $\mathcal{C}_{i-1}$, property (i’) implies that $\mathcal{C}_{i}$ also satisfies (i). Similarly (iii’) and the size condition on $\mathcal{C}_{i-1}$ implies $|\mathcal{C}_{i}|\leq\exp\left(\frac{r\log n\cdot\tau(n,m^{\prime})}{\epsilon}\right)^{i}$ as desired.
Let $t$ be such that $\mathcal{C}=\mathcal{C}_{t}$ (we implicitly show such a $t$ exists below). Note that (ii) holds by construction of $\mathcal{C}$, and (i) holds for $\mathcal{C}_{t}=\mathcal{C}$, so it so it remains to check (iii). Observe that by (ii’), we have for every $G\in\mathcal{C}_{i}$ that
$|G|\leq\max\left\{m,\left(1-\epsilon(\log n)^{-C}\right)^{i}\binom{n}{r}\right\}.$
Note that when $i\geq(\log n)^{\tilde{C}^{\prime}}$ for some sufficiently large constant $\tilde{C}^{\prime}$, we have
$\left(1-\epsilon(\log n)^{-C}\right)^{i}\binom{n}{r}<m.$
Thus we have $t\leq(\log n)^{\tilde{C}^{\prime}}$, and hence by the size condition
$|\mathcal{C}|=|\mathcal{C}_{t}|\leq\exp\left(\frac{r\log n\cdot\tau(n,m^{\prime})}{\epsilon}\right)^{(\log n)^{\tilde{C}^{\prime}}}\leq\exp\left((\log n)^{\tilde{C}}\tau(n,m)\right),$
where $\tilde{C}$ is some suitably large constant, completing the proof of the claim. ∎
Let $E=\max\left\{C^{\prime}pm,\tau(n,m)(\log n)^{C^{\prime}}\right\}$ where $C^{\prime}$ is a sufficiently large constant, and let $X_{E}$ denote the number of $F$-free $E$-edge subgraphs of $G^{r}_{n,p}$. By Claim 4.3 and linearity of expectation, we have
$\mathbb{E}[X_{E}]\leq|\mathcal{C}|\binom{m}{E}p^{E}\leq\exp\left((\log n)^{C^{\prime\prime\prime}}\tau(n,m)+\log\left(\frac{emp}{E}\right)\cdot E\right)\leq\exp(-E).$
Since $E\geq C_{2}pm\rightarrow\infty$ as $n\rightarrow\infty$, by Markov’s inequality, we have $\Pr[X_{E}\geq 1]\leq\mathbb{E}[X_{E}]\rightarrow 0$ as $n\rightarrow\infty$. Therefore, a.a.s.
$\mathrm{ex}(G^{r}_{n,p},F)\leq E=\max\left\{(\log n)^{C_{2}}\tau(n,m),~{}C_{2}pm\right\}.$
Combining our results here gives a criteria to establish tight bounds for random Turán problems.
Corollary 4.4.
Let $F$ be an $r$-graph with $\Delta(F)\geq 2$ such that $F$ is $(M,\gamma,\tau)$-balanced with $M(n)=\Theta(\mathrm{ex}(n,F))$, $\gamma=(\log n)^{\Theta(1)}$ and $\tau=n^{r-\frac{1}{d_{r}(F)}}(\log n)^{\Theta(1)}$. Then
$\mathrm{ex}(G^{r}_{n,p},F)=\left\{\begin{aligned} &\max\{\Theta(p\cdot\mathrm{ex}(n,F)),n^{r-\frac{1}{d_{r}(F)}}(\log n)^{\Theta(1)}\},~{}~{}&\text{if}~{}p\gg n^{-\frac{1}{d_{r}(F)}}\\ &(1+o(1))p\binom{n}{r},~{}~{}&\text{if}~{}n^{-r}\ll p\ll n^{-\frac{1}{d_{r}(F)}}.\\ \end{aligned}\right.$
Proof.
The upper bound comes from applying 1.5 with $m=M(n)$, and the lower bound comes from 4.1. ∎
We now prove 1.8 which we recall below.
Theorem 1.8.
Let $F$ be an $r_{0}$-graph with $2\leq\Delta(F)<|F|$ which is the spanning subgraph of a tight $r_{0}$-tree. If there exists a constant $C_{r_{0}}$ such that $F$ is $(M_{r_{0}},\gamma_{r_{0}},\tau_{r_{0}})$-balanced where $M_{r_{0}}(n)=C_{r_{0}}n^{r_{0}-1}$, $\gamma_{r_{0}}(n)=(\log n)^{C_{r_{0}}}$ and $\tau_{r_{0}}(n,m)=n^{r_{0}-\frac{1}{d_{r_{0}}(F)}}(\log n)^{C_{r_{0}}}$; then for all $r\geq r_{0}$, we have
$\mathrm{ex}(G^{r}_{n,p},F^{(r)})=\left\{\begin{aligned} &\max\{\Theta(pn^{r-1}),n^{r-\frac{1}{d_{r}(F^{(r)})}}(\log n)^{\Theta(1)}\},~{}~{}&\text{if}~{}p\gg n^{-\frac{1}{d_{r}(F^{(r)})}}\\ &(1+o(1))p\binom{n}{r},~{}~{}&\text{if}~{}n^{-r}\ll p\ll n^{-\frac{1}{d_{r}(F^{(r)})}}.\\ \end{aligned}\right.$
Proof.
By 3.4 (whose hypothesis is exactly the same as that of 1.8), we find for all $r\geq r_{0}$ that $F^{(r)}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced with $M(n)=\Theta(n^{r-1})$, $\gamma=(\log n)^{\Theta(1)}$ and $\tau=n^{r-\frac{1}{d_{r}(F)}}(\log n)^{\Theta(1)}$.
We claim that $\mathrm{ex}(n,F^{(r)})=\Theta(n^{r-1})$. Indeed, the lower bound comes from a star (i.e. every edge containing a given vertex $v\in[n]$), which is $F$-free since $\Delta(F)<|F|$ (meaning no vertex of $F^{(r)}$ is contained in every edge of $F^{(r)}$); and the upper bound comes from $M(n)=O(n^{r-1})$ and 3.4. We conclude the result by 4.4. ∎
5 Applying our Results
Here we apply the Lifting Theorems we have developed to give short proofs of Theorems 1.2 and 1.4 which give tight bounds for the random Turán problem for expansions of $K_{2,t}$ and theta graphs $\theta_{a,b}$. The details of the related result 1.3 for $K_{s,t}^{(r)}$ with general $s$ is deferred to an arXiv only appendix since the computations are cumbersome and since we prove stronger bounds for $s\geq 3$ in forthcoming work [16]. Our proof relies on 1.7, for which we need the following.
Lemma 5.1.
For all $a,b\geq 2$, the 3-graph $\theta^{(3)}_{a,b}$ is a spanning subgraph of a tight $3$-tree.
Proof.
By 3.3 it suffices to construct a tight $3$-tree whose 2-shadow contains a copy of $\theta_{a,b}$. Let $u$ and $v$ be the two end points of $\theta_{a,b}$, let $P_{i}$, $1\leq i\leq a$ be the $a$ paths, and let $w_{i,j}$, $11\leq j\leq b-$ 1, be the internal vertices of $P_{i}$, and for notational convenience we let $w_{i,0}=u$ and $w_{i,b}=v$ for all $i$. To build our tight tree, we begin by adding all of the edges of the form $\{w_{i,0},w_{i,1},w_{i,b}\}$ for all $i$, and iteratively given then we have added all edges of the form $\{w_{i,j-1},w_{i,j},w_{i,b}\}$ for some $21\leq j\leq b-$ 2 we add all edges of the form $\{w_{i,j},w_{i,j+1},w_{i,b}\}$. It is not difficult to see that this is a tight tree (since each edge we add adds the new vertex $w_{i,j+1}$) and that its 2-shadow contains the edge $w_{i,j},w_{i,j+1}$ for all $1\leq i\leq a$ and $11\leq j\leq b-$ 1 (with the $j=b-1$ case coming from the final edges $\{w_{i,b-2},w_{i,b-1},w_{i,b}\}$, and hence a copy of $\theta_{a,b}$. ∎
We now show that $\theta^{(r)}_{a,b}$ has balanced balanced supersaturation whenever $\theta_{a,b}$ does with $a\geq 3$.
Proposition 5.2.
Let $a\geq 3$ and $b\geq 2$ be such that there exists a constant $C_{2}>0$ such that $\theta_{a,b}$ is $(M_{2},C_{2},\tau_{2})$-balanced where
$M_{2}(n)=C_{2}n^{1+\frac{1}{b}}$
and
$\tau_{2}(n,m)=\max\left\{n^{1+\frac{a-1}{ab-1}},~{}m^{-\frac{1}{b-1}}n^{1+\frac{2}{b-1}}\right\}.$
Then for all $r\geq 3$, we have for all $0<p=p(n)\leq 1$ that a.a.s.
$\mathrm{ex}(G^{r}_{n,p},\theta_{a,b}^{(r)})=\left\{\begin{aligned} &\max\{\Theta(pn^{r-1}),n^{1+\frac{a-1}{ab-1}}(\log n)^{\Theta(1)}\},~{}~{}&\text{if}~{}p\gg n^{-r+1+\frac{a-1}{ab-1}}\\ &(1+o(1))p\binom{n}{r},~{}~{}&\text{if}~{}n^{-r}\ll p\ll n^{-r+1+\frac{a-1}{ab-1}}.\\ \end{aligned}\right.$
Proof.
We begin by using 1.7 to show that $\theta_{a,b}^{(3)}$ is balanced, and for this we need to check the hypothesis of 1.7. By Lemma 5.1, $\theta^{(3)}_{a,b}$ is a spanning subgraph of a tight 3-tree. Property (a) holds by hypothesis, and it is easy to check that $\gamma_{2}$ and $\tau_{2}$ satisfy properties (b) and (c). Thus Theorem 1.7 guarantees the existence of some sufficiently large constant $C_{3}$ such that the following holds. Let $M_{3}(n)=C_{3}n^{2}$, let $\gamma_{3}(n)=C_{2}C_{3}\log n$, and let
$A=A(n,m)=m^{\frac{ab-1}{2ab-a-1}}n^{-\frac{ab+a-2}{2ab-a-1}}.$
It is easy to check that property (d) in Theorem 1.7 holds, that is, for all sufficiently large $n$ and $m\geq M_{3}(n)=C_{3}n^{2}$ we have
$mn^{-2}\leq A(n,m)\leq\frac{m}{M_{2}(n)C_{3}\log n}=\frac{m}{C_{2}C_{3}n^{1+\frac{1}{b}}\log n}.$
Define the function
$\tau_{3}(n,m)=\max\left\{A(n,m)^{-2+\frac{a-1}{ab-1}}m,~{}\tau_{2}\left(n,\frac{m}{A(n,m)\cdot C_{3}\log n}\right)\cdot C_{3}\log n\right\}.$
By Theorem 1.7, we have that $\theta_{a,b}^{(3)}$ is $(M_{3},\gamma_{3},\tau_{3})$-balanced. Note that
$A(n,m)^{-2+\frac{a-1}{ab-1}}m=n^{1+\frac{a-1}{ab-1}},$
and we claim that
$\tau_{2}\left(n,\frac{m}{A(n,m)}\right)=\max\left\{n^{1+\frac{a-1}{ab-1}},~{}m^{-\frac{a}{2ab-a-1}}n^{1+\frac{3a}{2ab-a-1}}\right\}\leq n^{1+\frac{a-1}{ab-1}}.$
Indeed, using $m\leq C_{3}n^{2}$ and that $n$ is sufficiently large, it suffices to verify that $\frac{a-1}{ab-1}>\frac{a}{2ab-a-1}$, for which it suffices to show (using $a\geq 3$)
$(a-1)(2ab-a-1)-(ab-1)a$ $=(b-1)a^{2}-2ab+a+1$
$\geq(b-1)3a-2ab+a+1=(b-2)a+1>0$
which holds for $b\geq 2$. Thus, we have
$\tau_{3}(n,m)\leq n^{1+\frac{a-1}{ab-1}}(\log n)^{C_{3}}.$
Letting $\tau_{3}^{\prime}(n,m)=nn^{1+\frac{a-1}{ab-1}}(\log n)^{C_{3}}$, the above shows that $\theta^{(3)}_{a,b}$ is $(M_{3},\gamma_{3},\tau_{3}^{\prime})$-balanced and a spanning subgraph of a tight $3$-tree. The desired random Turán bound then follows for all $r\geq 3$ by 1.8. ∎
To utilize 5.2, we make use of the following known balanced supersaturation results (rephrased in the language of our paper).
Theorem 5.3 ( [10] Theorem 7.4).
For all $t\geq s\geq 2$, there exist sufficiently large $C_{2}>0$ such that $K_{s,t}$ is $(M_{2},C_{2},\tau_{2})$-balanced where
$M_{2}(n)=C_{2}n^{2-1/s}$
and
$\tau_{2}(n,m)=\max\left\{n^{\frac{2st-s-t}{st-1}},~{}m^{-s+1}n^{2s-1}\right\}.$
Theorem 5.4 ([8] Corollary 5.2).
For all $a\geq 100$ and $b\geq 3$, there exists a sufficiently large constant $C_{2}$ such that $\theta_{a,b}$ is $(M_{2},C_{2},\tau_{2})$-balanced where
$M_{2}(n)=C_{2}n^{1+\frac{1}{b}}$
and
$\tau_{2}(n,m)=\max\left\{n^{1+\frac{a-1}{ab-1}},~{}m^{-\frac{1}{b-1}}n^{1+\frac{2}{b-1}}\right\}.$
Theorems 1.2 and 1.4 now follow immediately from 5.2 together with these balanced supersaturation results (and the fact that $K_{2,t}=\theta_{t,2}$).
Appendix A Appendix
A.1 Comparing the Methods
Here we very informally show that 1.7, whose proof used both Greedy and Shadow Expansion, is essentially always at least as strong as 1.6, whose proof only used Shadow Expansion. For this we use the notation $O^{*}$ and $\Omega^{*}$ to denote asymptotic upper bounds and lower bounds that hold up to polylogarithmic factors, and throughout our argument we will be loose with logarithmic terms in order to make some of the calculations clearer.
Let $F$ be an $(r-1)$-graph which satisfies the hypothesis of 1.7 by being $(M_{r-1},\gamma_{r-1},\tau_{r-1})$-balanced, and let us make the extra mild assumption that $\gamma_{r-1}(n)=O^{*}(1)$. As we show in 3.1, this implies $\tau_{r-1}(n,m)=\Omega^{*}(n^{r-\frac{1}{d_{r}(F^{(r)})}})$. With this in mind, we consider applying 1.7 with $A=(m/n)^{1/(r-1)}$ for $m\geq\max\{M_{r-1}^{\frac{r-1}{r-2}}(n)n^{-\frac{1}{r-2}},n^{r-1}\}$ (i.e. for the range of $m$ that 1.6 is valid for). Note that $m\geq M_{r-1}^{\frac{r-1}{r-2}}(n)n^{-\frac{1}{r-2}}$ is equivalent to having $A=(m/n)^{1/(r-1)}\leq m/M_{r-1}(n)$, so (up to logarithmic factors) we have that the upper bound of property (d) in 1.7, and it is straightforward to verify the lower bound $A\geq mn^{1-r}$ for $r\geq 3$ and $m\geq n^{r-1}$, so we can indeed apply 1.7 with this value of $A$. We now observe that the first term in the maximum of $\tau_{r}(n,m)$ defined in 1.7 satisfies
$A^{-\frac{1}{d_{r}(F^{(r))}}}m\leq n^{r-\frac{1}{d_{r}(F^{(r)})}},$
with the inequality using $m\leq n^{r}$. This together with the claim above implies that up to logarithmic terms $\tau_{r}(n,m)$ equals $\tau_{r-1}(n,A^{-1}m)=\tau_{r-1}(n,n^{\frac{1}{r-1}}m^{\frac{r-2}{r-1}})$, and this is exactly what 1.6 gives.
A.2 Expansions of Complete Bipartite Graphs
Here we prove 1.3 giving tight bounds for random Turán numbers of $K_{s,t}$ expansions for $r$ sufficiently large. We need the following lemma.
Lemma A.1.
When $r\geq s+1$, $K^{(r)}_{s,t}$ is a spanning subgraph of a tight $r$-tree.
Proof.
By 3.3 it suffices to construct a tight $(s+1)$-tree whose 2-shadow contains a copy of $K_{s,t}$. Let $S$ and $T$ be the two parts of $K_{s,t}$ such that $|S|=s$ and $|T|=t$, where $S=\{v_{1},\dots,v_{s}\}$ and $T=\{w_{1},\dots,w_{t}\}$. Pick the $(s+1)$-edges $e_{i}=S\cup\{w_{i}\}$ for $1\leq i\leq t$. It is easy to check that $e_{1},\dots,e_{t}$ form a tight $(s+1)$-tree whose 2-shadow contains all $v_{i}w_{j}$, where $1\leq i\leq s$ and $1\leq j\leq t$. ∎
Next we prove balanced supersaturation results for $K^{(r)}_{s,t}$ using The balanced supersaturation results for $K_{s,t}$ proved by Morris and Saxton (5.3) together with the Shadow Expansion method (1.6) and the Greedy Expansion method (1.7).
Lemma A.2.
Let $t\geq s\geq 2$ and $r\geq s+1$. There exist sufficiently large $C_{r}$ such that for the functions
$M_{r}(n)=C_{r}n^{r-1},$
$\gamma_{r}(n)=(\log n)^{C_{r}},$
and
$\tau_{r}(n,m)=\max\left\{n^{\frac{2st-s-t}{st-1}},~{}m^{-\frac{(s-1)^{2}t}{2(st-1)(r-s)+(s-1)^{2}t}}n^{\frac{(s+1)((st-1)(r-s)+(s-1)^{2}t)}{2(st-1)(r-s)+(s-1)^{2}t}}\right\}(\log n)^{C_{r}},$
$K^{(r)}_{s,t}$ is $(M_{r},\gamma_{r}(n),\tau_{r}(n,m))$-balanced.
Proof.
We will first use 1.6 to obtain balanced supersaturation results for $K^{(s)}_{s,t}$, and then apply 1.7 to deal with $K^{(r)}_{s,t}$ when $r\geq s+1$.
Let $M_{2},C_{2},\tau_{2}$ be as in Theorem 5.3 and let $\gamma_{2}(n)$ be constantly $C_{2}$. In order to apply 1.6 with $F=K_{s,t}$ and $r=s$, we need to check the four properties (a), (b), (c) and (d). Property (a) of $F$ being $(M_{2},\gamma_{2},\tau_{2})$-balanced is immediate from Theorem 5.3, property (b) of having $M_{2}(n)n^{-\frac{s-2}{s-1}}=n^{1+\frac{1}{s(s-1)}}$ being non-increasing holds, and property (c) of $\gamma_{2}(n)$ being non-decreasing is immediate. For property (d), we observe that $\tau_{2}(n,m)$ is non-increasing with $m$ and that for every $s\geq 2$
$\tau_{2}(nx,mx^{\frac{s-2}{s-1}})=\max\left\{(nx)^{\frac{2st-s-t}{st-1}},~{}m^{-s+1}n^{2s-1}x^{s+1}\right\}$
is increasing with respect to $x$. So property (d) also holds. Hence by 1.6, there exists sufficiently large $C_{s}$ such that for the functions
$M_{s}(n)=\max\left\{M_{2}(n)^{s-1}n^{-s+2},~{}n^{s-1}\right\}(\log n)^{C_{s}}=n^{s-1+\frac{1}{s}}(\log n)^{C_{s}},$
$\gamma_{s}(n)=(\log n)^{C_{s}},$
and
$\tau_{s}(n,m)=\tau_{2}(n,n^{\frac{s-2}{s-1}}m^{\frac{1}{s-1}})(\log n)^{C_{s}}=\max\left\{n^{\frac{2st-s-t}{st-1}},~{}m^{-{1}}n^{s+1}\right\}(\log n)^{C_{s}},$
$K^{(s)}_{s,t}$ is $(M_{s},\gamma_{s},\tau_{s})$-balanced.
Note that this balanced supersaturation result for $K^{(s)}_{s,t}$ is almost the same as Lemma A.2 when $r=s$; the only difference is that $M_{s}(n)$ is slightly larger than expected. This forbids us from using induction on $r$ with $r=s$ as the base case. To cope with this obstacle, we will prove an inductive claim with a slightly weaker requirement on $M_{r-1}$. Precisely, it suffices to prove the following claim.
Claim A.3.
Let $t\geq s\geq 2$ and $r\geq s+1$. If there exist $C_{r-1}$ such that for the functions
$M_{r-1}(n)=\max\{C_{r-1}n^{r-2},n^{s-1+\frac{1}{s}}(\log n)^{C_{r-1}}\},$
$\gamma_{r-1}(n)=(\log n)^{C_{r-1}},$
and
$\tau_{r-1}(n,m)=\max\left\{n^{\frac{2st-s-t}{st-1}},~{}m^{-\frac{(s-1)^{2}t}{2(st-1)(r-1-s)+(s-1)^{2}t}}n^{\frac{(s+1)((st-1)(r-1-s)+(s-1)^{2}t)}{2(st-1)(r-1-s)+(s-1)^{2}t}}\right\}(\log n)^{C_{r-1}},$
$K^{(r-1)}_{s,t}$ is $(M_{r-1},\gamma_{r-1}(n),\tau_{r-1}(n,m))$-balanced, then there exist $C_{r}$ such that for the functions
$M_{r}(n)=C_{r}n^{r-1},$
$\gamma_{r}(n)=(\log n)^{C_{r}},$
and
$\tau_{r}(n,m)=\max\left\{n^{\frac{2st-s-t}{st-1}},~{}m^{-\frac{(s-1)^{2}t}{2(st-1)(r-s)+(s-1)^{2}t}}n^{\frac{(s+1)((st-1)(r-s)+(s-1)^{2}t)}{2(st-1)(r-s)+(s-1)^{2}t}}\right\}(\log n)^{C_{r}},$
$K^{(r)}_{s,t}$ is $(M_{r},\gamma_{r}(n),\tau_{r}(n,m))$-balanced.
Proof.
It is easy to check that $\gamma_{r-1}$ and $\tau_{r-1}$ satisfy properties (b) and (c) in Theorem 1.7. Thus by Lemma A.1, we can apply Theorem 1.7 with $F=K^{(r-1)}_{s,t}$, and we let $C_{r}$ be the constant given by this theorem. We let
$A(n,m)=m^{\frac{2(st-1)}{2(st-1)(r-s)+(s-1)^{2}t}}n^{-\frac{(s+1)(st-1)}{2(st-1)(r-s)+(s-1)^{2}t}}.$
We need to check property (d) in Theorem 1.7, that is, for all sufficiently large $n$ and $m\geq M_{r}(n)$,
$mn^{1-r}\leq A(n,m)\leq\frac{m}{M_{r-1}(n)\cdot C_{r}\log n}.$
The first inequality is equivalent to
$m^{1-\frac{2(st-1)}{2(st-1)(r-s)+(s-1)^{2}t}}\leq n^{r-1-\frac{(s+1)(st-1)}{2(st-1)(r-s)+(s-1)^{2}t}}.$
Since $m\leq n^{r}$, it suffices to have $1\leq\frac{(2r-s-1)(st-1)}{2(st-1)(r-s)+(s-1)^{2}t}=\frac{2(st-1)(r-s)+(s-1)(st-1)}{2(st-1)(r-s)+(s-1)^{2}t}$, which is equivalent to $t\geq 1$.
We still need to check
$A(n,m)\leq\frac{m}{M_{r-1}(n)\cdot C_{r}\log n},$
which is equivalent to
$M_{r-1}(n)\leq\frac{m^{1-\frac{2(st-1)}{2(st-1)(r-s)+(s-1)^{2}t}}n^{\frac{(s+1)(st-1)}{2(st-1)(r-s)+(s-1)^{2}t}}}{C_{r}\log n}.$
Note that $m\geq C_{r}n^{r-1}$, so it suffices to check the inequality above when $m=C_{r}n^{r-1}$, which for some appropriate $C^{\prime}$ is
$M_{r-1}(n)\leq{C^{\prime}n^{r-2+\frac{st+t+s-3}{2(st-1)(r-s)+(s-1)^{2}t}}}(\log n)^{-1}.$
Clearly, this holds when $r\geq s+2$. When $r=s+1$, we need
$\frac{1}{s}<\frac{st+t+s-3}{2(st-1)+(s-1)^{2}t},$
which is equivalent to
$0<(s-1)(s+t-2),$
which clearly holds.
Note that when omitting logarithmic factors, we have
$\tau_{r-1}\left(n,\frac{m}{A(n,m)}\right)$ $=\max\left\{n^{\frac{2st-s-t}{st-1}},\left(\frac{m}{A(n,m)}\right)^{-\frac{(s-1)^{2}t}{2(st-1)(r-1-s)+(s-1)^{2}t}}n^{\frac{(s+1)((st-1)(r-1-s)+(s-1)^{2}t)}{2(st-1)(r-1-s)+(s-1)^{2}t}}\right\}$ (22)
$=\max\left\{n^{\frac{2st-s-t}{st-1}},m^{x}n^{y}\right\},$
where
$x$ $=-\frac{(s-1)^{2}t}{2(st-1)(r-1-s)+(s-1)^{2}t}\cdot\left(1-\frac{2(st-1)}{2(st-1)(r-s)+(s-1)^{2}t}\right)$
$=-\frac{(s-1)^{2}t}{2(st-1)(r-s)+(s-1)^{2}t}$
and
$y$ $=-\frac{(s-1)^{2}t}{2(st-1)(r-1-s)+(s-1)^{2}t}\cdot\frac{(s+1)(st-1)}{2(st-1)(r-s)+(s-1)^{2}t}+\frac{(s+1)((st-1)(r-1-s)+(s-1)^{2}t)}{2(st-1)(r-1-s)+(s-1)^{2}t}$
$=\frac{s+1}{2(st-1)(r-1-s)+(s-1)^{2}t}\cdot\frac{(2(st-1)(r-1-s)+(s-1)^{2}t)((st-1)(r-s)+(s-1)^{2}t)}{2(st-1)(r-s)+(s-1)^{2}t}$
$=\frac{(s+1)((st-1)(r-s)+(s-1)^{2}t)}{2(st-1)(r-s)+(s-1)^{2}t)}.$
Also note that
$A(n,m)^{-\frac{1}{d_{r}\left(K^{(r)}_{s,t}\right)}}$ $=A(n,m)^{-r+s-\frac{(s-1)^{2}t}{st-1}}$ (23)
$=m^{-\frac{2(st-1)(r-s)+2(s-1)^{2}t}{2(st-1)(r-s)+(s-1)^{2}t}}n^{\frac{(s+1)((st-1)(r-s)+(s-1)^{2}t)}{2(st-1)(r-s)+(s-1)^{2}t}}$
$=m^{x}n^{y}.$
Recall that
$M_{r}(n)=C_{r}n^{r-1},$
$\gamma_{r}(n)=(\log n)^{C_{r}}\geq C_{r}\gamma_{r-1}(n)\log n,$
and
$\tau_{r}(n,m)$ $=\max\left\{n^{\frac{2st-s-t}{st-1}},~{}m^{-\frac{(s-1)^{2}t}{2(st-1)(r-s)+(s-1)^{2}t}}n^{\frac{(s+1)((st-1)(r-s)+(s-1)^{2}t)}{2(st-1)(r-s)+(s-1)^{2}t}}\right\}(\log n)^{C_{r}}$
$\geq\max\left\{A(n,m)^{-\frac{1}{d_{r}\left(K^{(r)}_{s,t}\right)}},\tau_{r-1}\left(n,\frac{m}{C_{r}A(n,m)\log n}\right)C_{r}\log n\right\}.$
The last inequality above comes from (23) and (22). Thus by Theorem 1.7 $K^{(r)}_{s,t}$ is $(M_{r},\gamma_{r},\tau_{r})$-balanced. ∎
We use Lemma A.2 together with 1.5 to prove Theorem 1.3.
Proof of Theorem 1.3.
By 4.1 it suffices to only prove the upper bound of Theorem 1.3 for the larger range of $p$. By Lemma A.2, we can apply Lemma 1.5 with $F=K^{(r)}_{s,t}$ which guarantees the existence of a sufficiently large constant $C_{1}$ such that for all $n\geq N_{r}$ and $m\geq C_{r}n^{r-1}$,
$\mathrm{ex}(K^{(r)}_{s,t},G^{r}_{n,p})\leq\max\left\{{C_{1}}pm,n^{\frac{2st-s-t}{st-1}}(\log n)^{C_{1}},~{}m^{-\frac{(s-1)^{2}t}{2(st-1)(r-s)+(s-1)^{2}t}}n^{\frac{(s+1)((st-1)(r-s)+(s-1)^{2}t)}{2(st-1)(r-s)+(s-1)^{2}t}}(\log n)^{C_{1}},\right\}.$ (24)
Let $M_{1}$ be such that
${C_{1}}pM_{1}=M_{1}^{-\frac{(s-1)^{2}t}{2(st-1)(r-s)+(s-1)^{2}t}}n^{\frac{(s+1)((st-1)(r-s)+(s-1)^{2}t)}{2(st-1)(r-s)+(s-1)^{2}t}}(\log n)^{C_{1}}.$ (25)
One can check that for some constant $C_{2}$
$M_{1}\leq p^{-\frac{2(st-1)(r-s)+(s-1)^{2}t}{2(st-1)(r-s)+2(s-1)^{2}t}}n^{\frac{s+1}{2}}(\log n)^{C_{2}}$ (26)
If $M_{1}\geq C_{r}n^{r-1}$, then we can apply (24), which by (25) and (26) gives
$\mathrm{ex}(K^{(r)}_{s,t},G^{r}_{n,p})$ $\leq\max\left\{n^{\frac{2st-s-t}{st-1}}(\log n)^{C_{1}},~{}C_{1}pM_{1}\right\}$ (27)
$\leq\max\left\{n^{\frac{2st-s-t}{st-1}}(\log n)^{C_{1}},~{}p^{\frac{(s-1)^{2}t}{2(st-1)(r-s)+2(s-1)^{2}t}}n^{\frac{s+1}{2}}(\log n)^{C_{2}}\right\}.$
If $M_{1}\leq C_{r}n^{r-1}$, let $m=C_{r}n^{r-1}$. Since $m\geq M_{1}$, by (25) we have
$m^{-\frac{(s-1)^{2}t}{2(st-1)(r-s)+(s-1)^{2}t}}n^{\frac{(s+1)((st-1)(r-s)+(s-1)^{2}t)}{2(st-1)(r-s)+(s-1)^{2}t}}(\log n)^{C_{1}}\leq{C_{1}}pm.$
Thus by (24),
$\mathrm{ex}(K^{(r)}_{s,t},G^{r}_{n,p})$ $\leq\max\left\{n^{\frac{2st-s-t}{st-1}}(\log n)^{C_{1}},~{}C_{1}pm\right\}$ (28)
$=\max\left\{n^{\frac{2st-s-t}{st-1}}(\log n)^{C_{1}},~{}C_{1}C_{r}pn^{r-1}\right\}.$
To summarize, by (27) and (28), for some sufficiently large constant $C_{3}$ we have
$\mathrm{ex}(K^{(r)}_{s,t},G^{r}_{n,p})\leq\max\left\{n^{\frac{2st-s-t}{st-1}}(\log n)^{C_{3}},~{}p^{\frac{(s-1)^{2}t}{2(st-1)(r-s)+2(s-1)^{2}t}}n^{\frac{s+1}{2}}(\log n)^{C_{3}},~{}C_{3}pn^{r-1}\right\}.$
From ease of analysis we ignore the logarithmic terms in these expressions. Let $\alpha=\alpha(r,s,t)>0$ be the unique value such that at $p=n^{-\alpha}$ the first and third terms equal each other and such that for smaller $p$ the first term is larger than the third, i.e. $\alpha=\frac{2st-s-t}{st-1}+1-r$. Similarly there exists some $\beta=\beta(r,s,t)>0$ such that at $p=n^{-\alpha}$ the last two terms are equal and such that for all smaller $p$ the second term is larger (this exists since the exponent for $p$ and $n$ in the third term is larger for $r\geq s+1$ and $s\geq 3$).
Observe that the second term will never achieve the maximum if $\beta\geq\alpha$, since in this case the second term will only beat the third term for $p=n^{-\gamma}$ with $\gamma\geq\beta$ but this value will be smaller than the second term at $p=n^{-\alpha}$ which is smaller than the third term which is smaller than the first term. A laborious computation shows that $\beta\geq\alpha$ holds precisely if
$r\geq s+1+\frac{s-1}{t-1}-\frac{s+t-2}{st-1}.$
This inequality holds for all $r\geq s+2$ since the fractional part is at most 1 due to $s\leq t$. At $r=s+1$ it holds provided $\frac{s-1}{t-1}-\frac{s+t-2}{st-1}\leq 0$, and in fact this holds for the even larger quantity $\frac{s-1}{t-1}-\frac{s+t-2}{st-1}+\frac{1}{(t-1)(st-1)}=\frac{t(s^{2}-2s+3-t)}{(t-1)(st-1)}$ provided $t\geq s^{2}-2s+3$ as desired. ∎